### v5

# appAppendix References

Abstract
As diffusion probabilistic models (DPMs) are being employed as mainstream models for Generative Artificial Intelligence (GenAI), the study of their memorization has attracted growing attention. Existing works in this field aim to establish an understanding of whether or to what extent DPMs learn via memorization. Such an understanding is crucial for identifying potential risks of data leakage and copyright infringement in diffusion models and, more importantly, for trustworthy application of GenAI. Existing works revealed that conditional DPMs are more prone to memorize training data than unconditional DPMs. And most data extraction methods developed so far target conditional DPMs. Although unconditional DPMs are less prone to data extraction, further investigation into these attacks remains essential since they serve as the foundation for conditional models like Stable Diffusion, and exploring these attacks will enhance our understanding of memorization in DPMs. In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) that leverages a time-dependent classifier trained on generated data as surrogate conditions to extract training data from unconditional DPMs. Empirical results demonstrate that it can extract training data in challenging scenarios where previous methods fail, and it is, on average, over 50% more effective across different scales of the CelebA dataset. Furthermore, we provide a theoretical understanding of memorization in both conditional and unconditional DPMs and why SIDE is effective.
1 Introduction
The diffusion probabilistic models (DPMs) [16, 30, 34] is a family of powerful generative models that learn the distribution of a dataset by first gradually destroying the structure of the data through an iterative forward diffusion process and then restoring the data structure via a reverse diffusion process. Due to their outstanding capability in capturing data distribution, DPMs have become the foundation models for many pioneering Generative Artificial Intelligence (GenAI) products such as Stable Diffusion [26], DALL-E 3 [2], and Sora [3]. Despite the widespread adoption of DPMs, a potential risk they face is data memorization, i.e., the risk of memorizing a certain proportion of the raw training samples. This could result in the generation of memorized (rather than new) samples via direct copying, causing data leakage, privacy breaches or copyright infringement, as highlighted in the literature [31, 32, 1, 11]. A current case argues that Stable Diffusion is a 21st-century collage tool, remixing the copyrighted creations of countless artists whose works were included in the training data[4]. Furthermore, data memorization also gives rise to data extraction attacks, which is one type of privacy attacks that attempt to extract the training data from a well-trained model. Notably, recent work by [5] demonstrated the feasibility of extracting training data samples from DPMs like Stable Diffusion [26], revealing the potential privacy and copyright dangers associated with them.
Several works have investigated the data memorization phenomenon in diffusion models. For example, it has been observed that there exists a strong correlation between training data memorization and conditional DPMs, i.e., conditional-trained DPMs are more prone to memorization [32]. Gu et al. [13] investigates influential factors on memorization behaviors via comprehensive experiments. They found that conditioning on random-labeled data can significantly trigger memorization, and unconditional models memorize much less training data. Though these works indicate that unconditional diffusion models are much less prone to data extraction attacks, investigating such attacks for unconditional diffusion models is still necessary. This is because 1) unconditional DPMs are the foundations of conditional generative models like Stable Diffusion, which adapts pre-trained unconditional DPMs as the essential part for conditional generation or combines them with conditional training approaches [26, 39]; 2) exploring data extraction attacks on unconditional DPMs enhances understanding of memorization mechanisms in both conditional and unconditional DPMs.
In this work, we propose a novel data extraction method named Surrogate condItional Data Extraction (SIDE) to extract training data from unconditional DPMs. This attack method uses a classifier trained on the same or similar training data as a surrogate condition for unconditional DPMs. SIDE can extract training data effectively by making the classifier time-dependent (to the diffusion sampling process). We empirically verify the effectiveness of SIDE on CIFAR-10 and different scales of the CelebA dataset (attack results in Figure 1). Additionally, to offer a theoretical analysis, we introduce a memorization metric to quantify the degree of memorization in DPMs by measuring the overlap between the generated and training data in a point-wise manner. Based on this metric, we explain why conditional DPMs are more likely to memorize training data, why random labelling can increase this memorization and why the SIDE method is effective for extracting training data in unconditional DPMs.
In summary, our main contributions are:
  • We propose a novel training data extraction method SIDE that leverages the implicit labels learned by a time-dependent classifier to extract training data from unconditional DPMs.
  • We introduce a novel metric to measure the degree of point-wise memorization in DPMs and present a theoretical analysis that explains 1) why conditional DPMs are more prone to memorization, 2) why random labels can lead to more memorization, and 3) why the SIDE method is effective for extracting training data in unconditional DPMs.
  • We evaluate the effectiveness of SIDE on CIFAR-10 and various scales of CelebA datasets and show that, on average, it can outperform the baseline method by more than 50%.
Figure 1: A few examples of the extracted images from a DDPM trained on a subset of the CelebA dataset using our SIDE method. Top: training images; bottom: extracted images.
2 Related Work
Diffusion Probabilistic Models  DPMs [30] such as Stable Diffusion [26], DALL-E 3 [2], Sora [3], Runway [26], and Imagen [28] have achieved state-of-the-art performance in image/video generation across a wide range of benchmarks [12]. These models can be viewed from two perspectives. The first is score matching [34], where diffusion models learn the gradient of the image distribution [35]. The second perspective involves denoising DPMs [16], which add Gaussian noise at various time steps to clean images and train models to denoise them. To conditionally sample from diffusion models, [12] utilizes a classifier to guide the denoising process at each sampling step. Additionally, [17] introduces classifier-free guidance for conditional data sampling using DPMs.
Memorization in Diffusion Models  Early exploration of memorization in large models was focused on language models [6, 19], which has motivated more in-depth research on DPMs [32, 13, 24]. Notably, a recent research on image-generation DPMs: [31] found that 0.5-2% of generated images duplicate training samples, a result concurrently reported by [5] in broader experiments for both conditional and unconditional diffusion models. Further studies [32] and [13] linked memorization to model conditioning, showing that conditional models are more prone to memorization. Several works have proposed methods to detect and mitigate memorization to address this issue. Wen et al. [38] proposed a method to detect whether a prompt can trigger memorization by inspecting the magnitude of text-conditional predictions. It delivers high accuracy even at the first generation step, with a single generation per prompt to detect whether a prompt can trigger memorization. Ren et al. [25] proposed two metrics based on findings on cross-attention in DPMs to detect memorization. To mitigate memorization, Chen et al.[7] proposed anti-memorization guidance to mitigate memorization during the sampling process. Instead of using guidance, Ren et al. [25] mitigated memorization by changing the attention score or masking out summary tokens in the cross-attention layer. Wen et al.[38] achieved this by minimizing the magnitude during inference. So far, the motivated data extraction attack or defend methods are primarily for conditional DPMs [5, 37], and studies have shown that extracting training data from unconditional DPMs can be much more challenging than that on conditional DPMs [13]. Though challenging, investigating data extraction methods in unconditional DPMs is still necessary. This is because 1) Unconditional DPMs are key parts in conditional models like Stable Diffusion, which adapt them for conditional generation, and 2) Exploring this approach enhances our understanding of memorization in DPMs. In response to these challenges, we propose a novel data extraction method called SIDE for unconditional DPMs and providing theoretical understanding of SIDE.
3 Proposed SIDE Method
3.1 Intuition
Previous studies [13, 32] have shown that conditional DPMs are more prone to memorization due to their reliance on explicit conditioning labels, such as class labels or text conditions, which we refer to as explicit informative labels or explicit labels. In contrast, unconditional DPMs do not have access to these explicit labels. Nevertheless, they still learn to form representation clusters from the training data, even though these clusters are not directly specified in the data itself [8]. We name these representation clusters as implicit informative labels or implicit labels. By formalizing the clustering patterns within the training data, it may be possible to construct implicit labels and make the unconditional DPMs conditional on these implicit labels to extract training data from unconditional DPMs.
In the following section, we will construct implicit informative labels for unconditional DPMs, convert the implicit labels into explicit ones, and then leverage the explicit labels to extract training data.
3.2 Constructing Implicit Informative Labels
Intuitively, one could use a classifier to generate (predict) implicit labels $y_{I}$ during the sampling process of the diffusion model. The classifier can be a normal classifier trained on the same dataset as the diffusion model. When such a classifier is not available, random labels or representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can also be used as the implicit labels. We assume an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process:
$\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\Big{(}\nabla_{\bm{x}}\log p_{\theta}^{t}({\bm{x}})+$
$\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\Big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (1)
where $x$ represents the state vector, $f(x,t)$ denotes the drift coefficient, $g(t)$ is the diffusion coefficient, $\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$ denotes the gradient of model $p_{\theta}$ given ${\bm{x}}$ at time $t$, and $\mathrm{d}w$ corresponds to the increment of the Wiener process.
We can use the classifier that generates the implicit labels to approximate the gradient in Eq. (3.2). However, a known challenge associated with neural network classifiers is their tendency towards miscalibration [14]. Specifically, the classifier could be overconfident or underconfident about its predictions. To mitigate the potential impact of miscalibration on the sampling process, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output on the diffusion path using power prior as follows:
$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right)$ (2)
Then, we have:
$\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\big{(}\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$
$+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w$ (3)
[…]
Figure 2: An illustration of our proposed Time-Dependent Knowledge Distillation (TDKD) method that trains a time-dependent classifier on a pseudo-labeled synthetic dataset.
3.3 Time-dependent Classifier
In Eq. (3.2), the classifier is denoted by $p_{\theta}^{t}(y|{\bm{x}})$ with $t$ being the timestep, implying its time-dependent nature. However, we do not have a time-dependent classifier but only a time-independent classifier. To address this problem, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier. The distillation process is illustrated in Figure 2. TDKD equips the classifier with time-dependent guidance during sampling. It operates in two steps: first, the network architecture is adjusted to accommodate time-dependent inputs, and the structure of the time-dependent module and modification are illustrated in Appendix C; second, a synthetic dataset and pseudo labels are created to distil knowledge from the normal classifier to its time-dependent counterpart.
As the original training dataset is unknown, we employ the target DPM to generate a synthetic dataset, following the generative data augmentation techniques [10, 9]. We then use the normal classifier trained on the original dataset to generate pseudo labels for the synthetic dataset. Finally, we train a time-dependent classifier on the labelled synthetic dataset. The objective of this training is defined as follows:
$\mathcal{L}_{distill}=D_{KL}\left(p_{\theta}(y_{I}|{\bm{x}}),p_{\theta}^{t}(y_{I}|{\bm{x}}_{t})\right)$ (4)
Overall Procedure of SIDE
With the trained time-dependent classifier $p_{\theta}^{t}(y|{\bm{x}}_{t})$ and the target DPM, our SIDE extracts training data from the model following a conditional generation process. Assume we condition on the label $y=c$. Firstly, we pick a set of values for $\lambda$, called $\mathcal{S}_{\lambda}$, to use in the SIDE attack. Then, we gather $N_{G}$ data samples for each value of $\lambda$ from the set. For each sampling timestep $t$, we compute the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ ($CE(\cdot)$ denotes the cross-entropy loss) and use the gradient and the target DPM to reverse the diffusion process. Next, we compute the similarity score for each generated image. Finally, we evaluate the attack performance using evaluation metrics. Defer to Appendix E for SIDE’s pseudocode.
4 Experiments
In this section, we first describe our experimental setup, introduce the performance metrics and then present the main evaluation results of our SIDE method. We empirically verify that the memorized images are not from the classifier. We also conduct an ablation study and hyper-parameter analysis to help understand the working mechanism of SIDE.
4.1 Experimental Setup
We evaluate the effectiveness of our method on three datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) [23] which consists of 5478 images, 2) a subset of the CelebA (CelebA-25000) [21] which contains 25,000 images and 2) CIFAR-10 containing 50,000 images. All the images are normalized to [-1,1]. We use the AdamW optimizer [22] with a learning rate of 1e-4 to train the time-dependent classifier. We train denoising DPMs with a discrete denoising scheduler (DDIM [33]) using the HuggingFace implementation [36]. All DPMs are trained with batch size 64. We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, respectively. We use ResNet34 [15] as the normal classifier.
4.2 Performance Metrics
Determining whether a generated image is a memorized copy of a training image is difficult, as $L_{p}$ distances in the pixel space are ineffective. Previous research addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection [31, 13]. However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization. Here, we propose two new memorization scores to solve these issues: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS).
The AMS and UMS metrics are defined as follows:
$AMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{1},\alpha,\beta\right)}}{N_{G}}$ (5)
$UMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)|}{N_{G}}$ (6)
where ${{\mathcal{D}}}_{1}$ is the generated dataset, ${\mathcal{D}}_{2}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring. $\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria. $\phi({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria. Mathematically, we can represent:
$\mathcal{F}(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\mathds{1}\Big{[}\max_{x_{j}\in{\mathcal{D}}_{2}}\gamma(x_{i},x_{j})\in[\alpha,\beta]\Big{]}$ (7)
$\phi(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{2},\gamma(x_{i},x_{j})\in[\alpha,\beta]\}$ (8)
where $\gamma$ represents the similarity/distance function. For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$ following [5], while for high-resolution datasets, we use the SSCD score as $\gamma$. In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity. (examples of different thresholds are shown in Figure 3) The thresholds for the normalized $L_{2}$ ditance [5] are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
The AMS averages the similarity scores across generated images, ensuring that memorized images are not overlooked. In contrast, the UMS quantifies distinct memorized instances by evaluating unique matches, thereby accounting for the uniqueness of the memorized images. By further categorizing the two scores into three levels—low, mid, and high—we obtain more comprehensive measurements for different types of memorization.
While [5, 7] introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is essential for assessing different types of copyright infringement, such as character or style copying [20, 27, 29]. Additionally, the UMS considers the number of generated images $N_{G}$, which was overlooked in [5]. The significance of $N_{G}$ lies in its non-linear impact on the UMS (see Appendix B for proof), indicating that UMS scores should not be compared across different values of $N_{G}$.
Figure 3: A comparison between the original training images (top row) and generated images (bottom row) by our SIDE method. The matches are classified into three categories: low similarity (SSCD score $<$ 0.5), mid similarity (SSCD score between 0.5 and 0.6), and high similarity (SSCD score $>$ 0.6). This classification highlights varying degrees of semantic resemblance among the image pairs.
Table 1: The AMS (%) and UMS (%) results at low, mid, and high levels. ‘Random’ denotes the baseline that generates images directly using the target unconditional DPM[5], while OL-TI is a variant of SIDE which is not trained using TDKD.
Dataset Method Low Similarity Mid Similarity High Similarity
AMS(%) UMS(%) AMS(%) UMS(%) AMS(%) UMS(%)
CelebA-HQ-FI Random 11.656 2.120 0.596 0.328 0.044 0.040
OL-TI 2.649 0.744 0.075 0.057 0.005 0.005
SIDE (Ours) 15.172 2.342 1.115 0.444 0.054 0.044
CelebA-25000 Random 5.000 4.240 0.100 0.100 0.000 0.000
OL-TI 0.164 0.152 0.000 0.000 0.000 0.000
SIDE (Ours) 8.756 6.940 0.224 0.212 0.012 0.012
CIFAR-10 Random 2.470 1.770 0.910 0.710 0.510 0.420
OL-TI 2.460 1.780 0.800 0.680 0.420 0.370
SIDE (Ours) 5.325 2.053 2.495 0.860 1.770 0.560
4.3 Main Results
We compare SIDE with the Random baseline proposed by Carlini et al.[5] and a variant of SIDE that substitutes the time-dependent classifier with a standard (time-independent) classifier. Here, ”TD” refers to the time-dependent classifier trained using our proposed TDKD method, ”TI” denotes the time-independent classifier, and ”OL” indicates training with original dataset labels.
The ”Random” baseline generates images directly using the target unconditional DPM, as described in [5]. We average the results across various values of $\lambda$ (defined in Eq. (3.2)), ranging from 5 to 9, with a detailed analysis provided in Section 4.3. It is important to note that $\lambda=0$ corresponds to the ”Random” baseline.
For each $\lambda$, including $\lambda=0$, we generate 50,000 images for CelebA and 10,000 images for CIFAR-10 to validate the proposed SIDE method. This effort results in one of the most enormous generated image datasets for studying the memorization of DPMs.
Table 2: This table presents the results of fitting a linear model to the relationship between the number of classes and both AMS and UMS.
Relationship Coefficient (×$\times 10^{-5}$) Intercept $R^{2}$ Correlation Coefficient
#Class vs. AMS 7.4 (positive) 0.115 0.637 0.80
#Class and UMS 6.1 (positive) 0.090 0.483 0.70
Figure 4: Hyper-parameter ($\lambda$) analysis on CelebA-HQ-FI. For high similarity, the best $\lambda$ for AMS and UMS are 16 and 13. For other similarity levels, the best $\lambda$ for AMS and UMS is 13.
Effectiveness of SIDE
The AMS and UMS results for the three datasets are presented in Table 1 (results of 95-th percentile of SSCD are listed in Appendix F). As shown, SIDE is highly effective in extracting memorized data across all similarity levels, particularly on the CelebA-25000 dataset, a task previously deemed unfeasible due to its large scale.
On CelebA-HQ-FI, SIDE increases mid-level AMS by 87% to 1.115% and UMS by 37% to 0.444%, with an average improvement of 20% across other levels. For the CelebA-25000 dataset, SIDE dramatically enhances AMS and UMS, achieving increases of 75% and 63% for low similarity and 124% and 112% for mid similarity. In the high similarity, SIDE excels at extracting memorized data.
On CIFAR-10, SIDE also outperforms the baselines across all similarity levels. For low similarity, it achieves an AMS of 5.325%, more than double Random’s 2.470%, and a UMS of 2.05% compared to 1.780%. For mid similarity, SIDE reaches 2.495% AMS and 0.860% UMS, significantly higher than Random’s 0.910% AMS and 0.710% UMS, respectively. In the high similarity category, SIDE achieves 1.770% AMS and 0.560% UMS, well ahead of Random’s 0.510% AMS and 0.420% UMS.
Effectiveness of TDKD
As shown in Table 1, time-independent classifiers perform significantly worse than their time-dependent counterparts on the two high-resolution CelebA datasets, achieving only about 10% of the effectiveness of classifiers trained using our TDKD method. This discrepancy arises because time-independent classifiers can provide accurate gradients only at the final timestep, whereas time-dependent classifiers deliver accurate gradients at each timestep. However, the performance gap narrows on the CIFAR-10 dataset due to its simplicity, consisting of only 10 classes. Thus, the gradients produced by the time-independent classifier are less prone to inaccuracy, which mitigates the limitations of the time-independent classifier.
Memorized Images Are Not From the Classifiers
There might be a concern that the extracted images may originate from the classifier rather than the target DPM. We argue that even if a time-dependent classifier could disclose its training images, these images would still originate from the target DPM because the classifier is distilled on the synthetic dataset generated by the target DPM. Furthermore, Table  1 demonstrates that employing a simple classifier reduces memorization compared to the baseline extraction method, indicating that the memorized images do not originate from the simple classifier as well.
Impact of the Classifier on SIDE
The classifier used to train SIDE is associated with a certain number of classes. Here, we conduct experiments to explore the relationship between the number of classes (of the classifier) and the extraction performance at a low similarity level, using 1,200 images per class with $\lambda=5$ on the CelebA-HQ-FI dataset. As shown in Table  2 and Figure 5, there exists a strong positive correlation: as the number of classes increases, both AMS and UMS improve. In summary, increasing number of classes positively affects both AMS and UMS, with a stronger impact on AMS. The UMS values reported here differ significantly from Table  1. This is because we only generated 1,200 images per class here, whereas 50,000 images per class in the previous experiment. Increasing the number of classes improves AMS more than UMS because a higher class count enables the classifier to better differentiate fine details, leading to more accurate matches at low similarity levels. UMS is less affected since it relies more on the diversity across images, which is constrained by the smaller number of images per class in this experiment.
Figure 5: Scatter plots showing the relationship between the number of classes and two performance metrics, AMS (left) and UMS (right). The fitted regression lines demonstrate a positive correlation in both cases.
Hyper-parameter Analysis
Here, we test the sensitivity of SIDE to its hyper-parameter $\lambda$. We generate 50,000 images for each integer value of $\lambda$ within the range of [0, 50]. As shown in Figure 4, the memorization score increases at first, reaching its highest, then decreases as $\lambda$ increases. This can be understood from sampling SDE Eq. (3.2). Starting from 0, the diffusion models are unconditional. As $\lambda$ increases, the diffusion models become conditional, and according to Theorem 1, the memorization effect will be triggered. However, when $\lambda$ is too large, the generated images will overfit the classifier’s decision boundaries, leading to a low diversity and ignoring the data distribution. Consequently, the memorization score decreases.
5 Theoretical Understanding for Memorization in Conditional Generative Models
In this section, we introduce a metric to quantify the memorization capacity of generative models and provide a theoretical explanation for the data memorization in conditional diffusion models and why SIDE is effective.
5.1 Memorization Metric
Intuitively, the memorization of fixed training data points (i.e., point-wise memorization) can be quantified by the degree of overlap between the generated distribution and the distributions centered at each training data point. Given a generative model and training dataset, we propose the following memorization metric to quantify the memorization of the training data points in the model.
Definition 1 (Point-wise Memorization).
Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset ${\mathcal{D}}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, degree of memorization in $p_{\theta}$ of ${\mathcal{D}}$ is defined as:
${\mathcal{M}}_{p}({\mathcal{D}};p_{\theta})=\frac{1}{|{\mathcal{D}}|}\sum_{{\bm{x}}_{i}\in{\mathcal{D}}}{\int{p_{\theta}}}({\bm{x}})\log\frac{p_{\theta}({\bm{x}})}{q({\bm{x}},{\bm{x}}_{i},\epsilon)}\,d{\bm{x}}$ (9)
where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ denotes the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the probability distribution centered at training sample ${\bm{x}}_{i}$ within a small radius $\epsilon$.
A straightforward choice for $q({\bm{x}},{\bm{x}}_{i},\epsilon)$ is the Dirac delta function centered at training data point ${\bm{x}}_{i}$: $q({\bm{x}},{\bm{x}}_{i},\epsilon)=\delta({\bm{x}}-{\bm{x}}_{i})$. However, this would make Eq. (9) uncomputable as the Dirac delta function is zero beyond the $\epsilon$-neighborhood. Alternatively, we could let $q({\bm{x}},{\bm{x}}_{i},\epsilon)\sim{\mathcal{N}}(x_{i};\epsilon I)$. Note that in Eq. (9), a smaller value of ${\mathcal{M}}_{p}({\mathcal{D}};\theta)$ indicates more overlap between the two distributions and thus more memorization. As $\epsilon$ tends to 00, the measured memorization becomes more accurate, with the limit characterizing the memorization capability of $p_{\theta}$.
5.2 Theoretical Analysis
Based on the point-wise memorization metric, we perform a theoretical understanding that explains why conditional generative models achieve better memorization capacity. Our analysis is based on the concept of informative labels, which refers to information that can partition a dataset into various disjoint subsets. And we demonstrate that provided with an informative label, the conditional generative models achieve stronger memorization capacity.
Informative Labels
The concept of informative labels has been previously discussed as class labels [13]. In this work, we introduce a more general definition that encompasses both class labels and random labels as special cases. We formally define an informative label as follows.
Definition 2 (Informative Label).
Let ${\mathcal{Y}}$ be an attribute of the data taking values in $\{y_{i}\}_{i=1}^{C}$. We refer to ${\mathcal{Y}}$ as an informative label if we can partition the dataset into several mutually disjoint subsets $\{{\mathcal{D}}_{i}\}_{i=1}^{C}$ based on the values of the attribute ${\mathcal{Y}}$.
Here, the informative labels are not limited to the conventional class labels; they can also be text captions, features, or cluster information that can be used to group the training samples into subsets. The above definition states that an information label should be able to differentiate a subset of samples from others. An extreme case is that all samples have the same label; in this case, the label is not informative. According to our definition, class-wise and random labels are special cases of informative labels. Informative labels can be either explicit like class/random labels and text captions, or implicit like salient features or clusters.
Main Results
Here we introduce our main theorem on the memorization mechanism of conditional generative models and explain why SIDE is effective. We use ${\mathcal{D}}_{i}$ to denote the subset consisting of data with informative label ${\mathcal{Y}}=y_{i}$. In addition, we denote the distribution of data in the original dataset ${\mathcal{D}}$ by $p$. The distribution has expectation $\mu$ and covariance matrix $\Sigma$. Similarly, for each attribute $y_{i}$, we use $p_{i}$ to denote the distribution of the data with this attribute, and use $\mu_{i}$ and $\Sigma_{i}$ to denote its expectation and covariance matrix.
Theorem 1.
If a generative model $p_{\theta_{i}}$ matches the target distribution $p_{i}$ well for each $i$, i.e. $TV(p_{i},p_{\theta_{i}})=0$, and the data distribution for some informative label $y_{c}$ satisfies $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$, we have with probability 1:
$\lim_{|{\mathcal{D}}_{c}|\to\infty}\lim_{\epsilon\to 0}\frac{{\mathcal{M}}_{p}({\mathcal{D}}_{c},p_{\theta_{c}})}{{\mathcal{M}}_{p}({\mathcal{D}}_{c},p_{\theta})}\leq 1$ (10)
where $p_{\theta_{i}}$ denotes the distribution of generated data conditioned on ${\mathcal{Y}}=y_{i}$.
The proof can be found in Appendix A. Theorem 1 states that any form of information labels can enhance memorization capacity in conditional generative models, including conventional class labels and random labels. The condition $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$ is reasonable because it ensures that the spread (variance) of the data within each class is not excessively large.
$\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{c}}{\left({\bm{x}}_{i}-\mu_{c}\right)^{T}\left({\bm{x}}_{i}-\mu_{c}\right)}}{|{\mathcal{D}}_{c}|}\leq||\mu-\mu_{c}||_{2}^{2}$ (11)
To understand this, we look at the derivation Eq. (11) from $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$. For a well-separated class, this squared distance should provide an upper bound on how spread out the class data can be. If $tr(\Sigma_{c})$ were large relative to $||\mu-\mu_{c}||_{2}^{2}$, it would imply that data samples represented by informative label $c$ are highly spread out, making it harder for the class labelled with $c$ to be well-formed and distinct. Thus, the condition ensures that the data samples form tight clusters around their means, making it reasonable for the class to be well-defined and easily detachable.
This also explains why SIDE works. Chen et al. [8] found that unconditional DPMs learn clustered representations from training data. The data samples within each class should not be widely dispersed for effective clustering. Then, the samples should satisfy Eq. 11. So, although these clustered representations are implicit, we could extract training data from unconditional DPMs if we can convert them from implicit informative labels into explicit labels. When conditioning on the explicit labels, the DPMs will exhibit more memorization effects indicated by Theorem 1, which makes room for data extraction attacks.
6 Limitations:
SIDE has certain limitations that warrant further research. SIDE mainly uses human-labelled data to extract data, which is scarce and limits its practical use. Future work could investigate hybrid approaches for generating implicit labels, such as employing unsupervised clustering or utilizing pretrained models to label synthetic datasets.
7 Conclusion
In this paper, we distinguish the difference between explicit labels and implicit labels and propose a novel method, SIDE, to extract training data from unconditional DPMs by constructing surrogate conditions. The core of SIDE lies in training a time-dependent classifier using TDKD technique. We empirically validate effectiveness of SIDE on subsets of CelebA and CIFAR-10 datasets with two new memorization scores: AMS and UMS. In addition, we introduce a point-wise memorization metric to quantify memorization effects in DPMs for theoretical analysis. Finally, we introduce theoretical understanding of memorization in conditional generative models, offering a generalized definition of informative labels, clarifying random labels can also be informative and explaining why SIDE works. Our work aims to deepen understanding of memorization in DPMs and inspire further methods to mitigate memorization.
Supplementary Material
Broader Impacts
Our work introduces a memorization metric that not only quantifies memorization effects in diffusion models but also extends to deep learning models more broadly. This contribution is significant in enhancing our understanding of when and how models memorize training data, which is critical for addressing concerns about data privacy and model robustness. By providing a theoretical framework for conditional memorization, we pave the way for developing effective memorization mitigation strategies tailored for diffusion models. These advancements can lead to the design of more secure and trustworthy AI systems, reducing the risk of potential data leakage while fostering greater accountability in the use of generative technologies. Ultimately, our findings aim to empower researchers and practitioners to create models that better respect privacy, thus benefiting the wider AI community.
Appendix A Proof of Theorem1
By the assumption $TV(p_{i},p_{\theta_{i}})=0$ in the theorem, we can replace $p_{\theta_{c}}$ with $p_{c}$ and $p_{\theta}$ with $p$ to simplify the notation without compromising correctness. Firstly, we compute $\mathcal{M}_{point}(\mathcal{D}_{c};p_{c})$ here:
[…]
In the following, we will omit the expression ”almost surely”. Similarly, for $\mathcal{M}_{p}(\mathcal{D}_{c};p)$:
$\begin{split}\mathcal{M}_{p}(\mathcal{D}_{c};p)&=o(\frac{1}{\epsilon})+\frac{1}{2\epsilon|\mathcal{D}_{c}|}\sum_{i=1}^{|\mathcal{D}_{c}|}(tr(\Sigma)+(\mu-z_{i})^{T}(\mu-z_{i}))\\ &\rightarrow o(\frac{1}{\epsilon})+\frac{1}{2\epsilon}(tr(\Sigma)+tr(\Sigma_{c})+||\mu-\mu_{c}||^{2}_{2})\end{split}$ (13)
By definition, $p(x)=\sum_{k}\pi_{k}p_{k}(x)$ ,where $\pi_{k}$ denotes the probability that the informative label $\mathcal{Y}$ equals $y_{k}$.
$\begin{split}\Sigma=\underset{x\sim p(x)}{\mathrm{Cov}}(x)&=\mathbb{E}_{p}[xx^{T}]-\mathbb{E}_{p}[x]\mathbb{E}_{p}[x]^{T}\\ &=\sum_{k}\pi_{k}\mathbb{E}_{p_{k}}[zz^{T}]-\mu\mu^{T}\\ &=\sum_{k}\pi_{k}(\mu_{k}\mu_{k}^{T}+\Sigma_{k})-\mu\mu^{T}\\ &=\mathbb{E}_{\pi}[\Sigma_{k}]+\mathbb{E}_{\pi}[(\mu-\mu_{k})(\mu-\mu_{k})^{T}]\end{split}$ (14)
Finally, we have
$\begin{split}\lim_{|\mathcal{D}|_{c}\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{\mathcal{M}_{p}(\mathcal{D}_{c};p_{c})}{\mathcal{M}_{p}(\mathcal{D}_{c};p)}&=\frac{2tr(\Sigma_{c})}{tr(\Sigma_{c})+\mathbb{E}_{\pi}[tr(\Sigma_{k})]+\mathbb{E}_{\pi}[||\mu-\mu_{k}||^{2}_{2}]+||\mu-\mu_{c}||^{2}_{2}}\end{split}$ (15)
Here we use $\mathbb{E}_{\pi}[tr(\Sigma_{k})]$ and $\mathbb{E}_{\pi}[||\mu-\mu_{k}||_{2}^{2}]$ to denote $\sum_{k}\pi_{k}tr(\Sigma_{k})$ and $\sum_{k}\pi_{k}||\mu-\mu_{k}||_{2}^{2}$.
Clearly, by Jensen’s inequality we have:
$\mathbb{E}_{\pi}[\lim_{|\mathcal{D}_{c}|\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{\mathcal{M}_{p}(\mathcal{D}_{c};p_{c})}{\mathcal{M}_{p}(\mathcal{D}_{c};p)}]\leq 1$ (16)
The equality holds only when $\mu_{k}$ and $tr(\Sigma_{k})$ are the same for all $k$. And under the assumption $tr(\Sigma_{c})\leq||\mu-\mu_{c}||_{2}^{2}$, we have:
$\lim_{|\mathcal{D}_{c}|\rightarrow\infty}\lim_{\epsilon\rightarrow 0}\frac{\mathcal{M}_{p}(\mathcal{D}_{c};p_{c})}{\mathcal{M}_{p}(\mathcal{D}_{c};p)}\leq 1$ (17)
Hence we have completed the proof of Theorem1. ∎
Appendix B Explanation UMS’s importance in fair comparison
By highlighting the significance of $N_{G}$, UMS allows a fair comparison among extraction methods. We explain this both theoretically and experimentally. Let $N_{\mathrm{umem}}$ denote the number of uniquely memorized images, $M$ the dataset size, and $p_{\gamma}(i)$ the probability that image $i$ is memorized. To find $\mathbb{E}(N_{\mathrm{umem}})$, we introduce a new variable $I_{i}$, representing the generation of image $i$ in $N_{G}$ trials:
$p(I_{i})=1-\left(1-p_{\gamma}(i)\right)^{N_{G}}$ (18)
Using the linearity of expectation, we derive: $\mathbb{E}\left(N_{\mathrm{mem}}\right)$
$\mathbb{E}\left(N_{\mathrm{umem}}\right)$ (19)
$=\mathbb{E}\left(I_{1}\right)+\mathbb{E}\left(I_{2}\right)+\cdots+\mathbb{E}\left(I_{M}\right)$ (20)
$=\sum_{i=1}^{M}{1-\left(1-p_{\gamma}\left(i\right)\right)^{N_{G}}}$ (21)
The importance of $N_{G}$ lies in its impact on the non-linear expectation of uniquely memorized images. Comparing UMS across different $N_{G}$ values is flawed because varying $N_{G}$ leads to different outcomes, underscoring the need for consistent $N_{G}$ values to ensure fair comparisons. In [5], the significance of $N_{G}$ was overlooked, as they only reported the number of uniquely extracted images. Our theoretical analysis accurately aligns with behavior observed in the experimental data in Fig 6.
Appendix C Refinement ResNet block (Fig. 7)
Figure 6: Validation of $N_{G}$ ’s significance
Figure 7: Refinement ResNet block with time-dependent module integration. This block diagram depicts the insertion of a time module within a conventional ResNet block architecture, allowing the network to respond to the data’s timesteps. Image $x_{BN}$ is the image processed after the first Batch Normalization Layer.
The integration of the time module directly after batch normalization within the network architecture is a reasonable design choice rooted in the functionality of batch normalization itself. Batch normalization standardizes the inputs to the network layer, stabilizing the learning process by reducing internal covariate shifts. By positioning the time module immediately after this normalization process, the model can introduce time-dependent adaptations to the already stabilized features. This placement ensures that the temporal adjustments are applied to a normalized feature space, thereby enhancing the model’s ability to learn temporal dynamics effectively.
Moreover, the inclusion of the time module at a singular point within the network strikes a balance between model complexity and temporal adaptability. This singular addition avoids the potential redundancy and computational overhead that might arise from multiple time modules. It allows the network to maintain a streamlined architecture while still gaining the necessary capacity to handle time-varying inputs.
Appendix D Results on Classifier Choice
The analysis of classifier performance across varying numbers of classes reveals interesting patterns for both AMS and UMS. As shown in Figure 5, the scatter plots highlight a positive relationship between the number of classes and the performance metrics. For AMS, the regression line suggests a stronger relationship (R² = 0.637), implying that as the number of classes increases, the AMS metric improves with a moderately strong association. In contrast, the relationship between the number of classes and UMS, while still positive, exhibits a slightly weaker connection (R² = 0.483).
These results suggest that classifier performance, particularly as measured by AMS, benefits more significantly from an increase in the number of classes compared to UMS. The shaded regions in the plots represent the 95% confidence intervals, indicating the range of uncertainty around the fitted regression lines. Overall, the findings imply that the choice of classifier could have a notable impact on AMS , with a less pronounced but still meaningful effect on UMS.
Appendix E Pseudocode for SIDE method
The following Algorithm 1 outlines the detailed steps of the SIDE method for extracting training data from unconditional diffusion models. This algorithm combines the construction of implicit informative labels, the training of a time-dependent classifier, and the conditional generation process to extract valuable training samples effectively.
Algorithm 1 SIDE Method for Extracting Training Data from Unconditional Diffusion Models
1:Input:
  • Unconditional Diffusion Probabilistic Model (DPM) $p_{\theta}(\bm{x})=\mathcal{N}(\bm{\mu},\Sigma_{p_{\theta}})$
  • Pre-trained Classifier $p_{\theta}(y_{I}\mid\bm{x})$
  • Hyperparameter set $\mathcal{S}_{\lambda}$
  • Number of generated samples per $\lambda$, $N_{G}$
  • Target label $y=c$
2:Output: Extracted training data $\mathcal{D}_{extracted}$
3:Construct Implicit Informative Labels
4:Input: Unconditional DPM $p_{\theta}(\bm{x})$, Classifier $p_{\theta}(y_{I}\mid\bm{x})$
5:Define sampling process with implicit labels:
$\mathrm{d}\bm{x}$ $=\left[f(\bm{x},t)-g(t)^{2}\left(\nabla_{\bm{x}}\log p_{\theta}^{t}(\bm{x})+\lambda\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}\mid\bm{x})\right)\right]\mathrm{d}t+g(t)\mathrm{d}\bm{w}$ (22)
6:Calibrate Classifier Output
7:Adjust classifier probabilities using power prior:
$p^{t}_{\theta}\left(\bm{x}\mid y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}\mid\bm{x}\right)p^{t}_{\theta}\left(\bm{x}\right)$ (23)
8:Train Time-Dependent Classifier via TDKD
9:Input: Pre-trained Classifier $p_{\theta}(y_{I}\mid\bm{x})$, Synthetic Dataset $\mathcal{D}_{synthetic}$
10:Output: Time-Dependent Classifier $p_{\theta}^{t}(y_{I}\mid\bm{x}_{t})$
11:Initialize Time-Dependent Classifier architecture with time-dependent modules
12:Generate synthetic dataset using DPM:
$\mathcal{D}_{synthetic}=\{\bm{x}^{(i)}\}_{i=1}^{N_{synthetic}}\sim p_{\theta}(\bm{x})$ (24)
13:Generate pseudo labels using pre-trained classifier:
$y_{I}^{(i)}=p_{\theta}(y_{I}\mid\bm{x}^{(i)})$ (25)
14:Train Time-Dependent Classifier by minimizing KL divergence:
$\mathcal{L}_{distil}=D_{KL}\left(p_{\theta}(y_{I}\mid\bm{x}^{(i)})\,\|\,p_{\theta}^{t}(y_{I}\mid\bm{x}_{t}^{(i)})\right)$ (26)
15:Overall SIDE Procedure
16:Input: Trained Time-Dependent Classifier $p_{\theta}^{t}(y_{I}\mid\bm{x}_{t})$, Target Label $y=c$, Hyperparameter set $\mathcal{S}_{\lambda}$, Number of samples $N_{G}$
17:Initialize empty dataset $\mathcal{D}_{extracted}$
18:for each $\lambda\in\mathcal{S}_{\lambda}$ do
19:     for $i=1$ to $N_{G}$ do
20:         Sample $\bm{x}_{0}$ conditioned on $y=c$ using Eq. (22)
21:         Compute gradient of cross-entropy loss:
$\nabla_{\bm{x}_{t}}\mathcal{L}_{CE}(c,p_{\theta}^{t}(y\mid\bm{x}_{t}))$ (27)
22:         Reverse diffusion process using computed gradient
23:         Generate similarity score for $\bm{x}_{t}$
24:         Append $\bm{x}_{t}$ to $\mathcal{D}_{extracted}$ with similarity score
25:     end for
26:end for
27:Evaluate Attack Performance
28:Compute evaluation metrics on $\mathcal{D}_{extracted}$
29:return $\mathcal{D}_{extracted}$
Appendix F SSCD-related metrics
In our evaluation of the proposed method, it is essential to include a discussion of the 95th percentile SSCD metric alongside our newly introduced metrics, AMS and UMS. While 95th percentile SSCD metric has its limitations, it still serves as a useful reference point for assessing the relative similarity between extracted images and the original training dataset.
As shown in Table 3, we provide a comprehensive comparison of performance across different datasets.
In addition, we report detailed metrics in the context of our experiments on the CelebA-HQ-FI dataset, further illustrating our approach’s effectiveness.
Table 3: Performance Comparison on CelebA-HQ-FI and CelebA-25000 Datasets
Dataset Method Top 0.01% Top 0.05% Top 0.1% Top 0.5% Top 1.0% Top 5.0%
CelebA-HQ-FI Uncond 0.636 0.604 0.584 0.524 0.498 0.433
SIDE(ours) 0.680 0.639 0.618 0.567 0.544 0.485
CelebA-25000 Uncond 0.545 0.518 0.505 0.471 0.455 0.404
SIDE(ours) 0.585 0.554 0.539 0.506 0.491 0.450
Appendix G Extended SIDE for Text-to-Image Diffusion Models
We extend the SIDE method for a text-to-image diffusion model in a fine-tuning scenario. Instead of using a classifier as a surrogate condition to extract training data, we use random strings as surrogate conditions. The attack setting is simple and practical: the attacker can access the fine-tuned model and modify the text inputs in the text-to-image pairs. Notably, the attacker does not have access to the training data. We name this method as Extended SIDE.
G.1 Attack Setting
The attack setting for the Extended SIDE method involves two primary components: the victim’s fine-tuning process and the attacker’s access to the fine-tuned model. Below, we provide a detailed description of each component.
  • Victim Fine-Tuning:
    The victim fine-tunes a text-to-image diffusion model, denoted as $p_{\theta}(\bm{x}\mid s)$, using a dataset of image-text pairs $\mathcal{D}_{train}=\{(s_{i},\bm{x}_{i})\}_{i=1}^{N}$. In this context:
    • $s_{i}\in\mathcal{S}_{s}$ are random strings crafted by the attacker to serve as surrogate text prompts.
    • $\bm{x}_{i}\in\mathbb{R}^{d}$ represents the corresponding images associated with each random string.
    The fine-tuning objective is to optimize the model parameters $\theta$ to minimize the loss function over the training data:
    $\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(s,\bm{x})\sim\mathcal{D}_{train}}\left[\mathcal{L}(p_{\theta}(\bm{x}\mid s),\bm{x})\right]$ (28)
    where $\mathcal{L}$ denotes the loss function, such as mean squared error (MSE) or cross-entropy loss. The use of random strings as surrogate conditions aims to induce memorization within the model, causing it to overfit specific image-text associations during the fine-tuning process.
  • Attacker’s Access:
    After the fine-tuning phase, the attacker gains access to the fine-tuned model $p_{\theta}(\bm{x}\mid s)$. This access allows the attacker to perform the following actions:
    • Input Manipulation: The attacker can input any text prompt $s\in\mathcal{S}_{s}$ into the model to generate corresponding images:
      $\bm{x}\sim p_{\theta}(\bm{x}\mid s)$ (29)
    • Data Extraction: By systematically querying the model with the same set of random strings $\mathcal{S}_{s}$, the attacker can generate multiple images $\{\bm{x}_{i,j}\}_{j=1}^{N_{G}}$ for each string $s_{i}$. The goal is to identify images that closely resemble the memorized training samples:
      $\bm{x}_{i,j}\sim p_{\theta}(\bm{x}\mid s_{i}),\quad j=1,\dots,N_{G}$ (30)
    • Memorization Detection: The attacker analyzes the consistency and similarity among the generated images for each random string $s_{i}$. A low variance $\sigma_{s_{i}}^{2}$ among the samples indicates potential memorization:
      $\sigma_{s_{i}}^{2}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i,j}-\overline{\bm{x}}_{s_{i}}\|^{2}$ (31)
      where $\overline{\bm{x}}_{s_{i}}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i,j}$ is the mean image for string $s_{i}$. If $\sigma_{s_{i}}^{2}$ falls below a predefined threshold $\tau$, the attacker considers $\overline{\bm{x}}_{s_{i}}$ as a potential memorized training sample.
  • Assumptions:
    • Knowledge of Surrogate Strings: The attacker is aware of the set of random strings $\mathcal{S}_{s}$ used during the victim’s fine-tuning process.
    • Model Query Capability: The attacker can perform a large number of queries to the fine-tuned model with each random string $s_{i}$ to generate sufficient samples for variance analysis.
    • No Access to Training Data: The attacker does not have direct access to the victim’s training dataset $\mathcal{D}_{train}$.
  • Attack Goals:
    • Data Reconstruction: To reconstruct or approximate the training images $\bm{x}_{i}$ by exploiting the model’s memorization of the random string-image associations.
    • Privacy Breach: To demonstrate the potential privacy risks associated with fine-tuned diffusion models by extracting sensitive or proprietary training data.
G.2 Extended SIDE Method
The Extended SIDE method leverages random strings as surrogate conditions to directly extract training images from a fine-tuned diffusion model. Unlike the original SIDE method, which relies on a classifier to guide the extraction process, this approach simplifies the attack by using random strings to induce and exploit memorization within the model.
Algorithm 2 Extended SIDE Method for Extracting Training Data Using Random Strings
1:Input:
  • Fine-tuned diffusion model $p_{\theta}(\bm{x}\mid s)$
  • Set of random strings $\mathcal{S}_{s}$
  • Number of samples per string $N_{G}$
  • Variance threshold $\tau$
2:Output: Extracted data $\mathcal{D}_{\text{extracted}}$
3:Initialize $\mathcal{D}_{\text{extracted}}\leftarrow\emptyset$
4:for each $s\in\mathcal{S}_{s}$ do
5:     Generate samples $\{\bm{x}_{i}^{(j)}\}_{j=1}^{N_{G}}$ where $\bm{x}_{i}^{(j)}\sim p_{\theta}(\bm{x}\mid s)$
6:     Compute mean image:
$\overline{\bm{x}}_{s}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i}^{(j)}$ (32)
7:     Compute sample variance:
$\sigma_{s}^{2}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i}^{(j)}-\overline{\bm{x}}_{s}\|^{2}$ (33)
8:     if $\sigma_{s}^{2}<\tau$ then
9:         $\mathcal{D}_{\text{extracted}}\leftarrow\mathcal{D}_{\text{extracted}}\cup\{\overline{\bm{x}}_{s}\}$
10:     end if
11:end for
12:return $\mathcal{D}_{\text{extracted}}$
The Extended SIDE method operates on the principle that random strings, when used as surrogate conditions during fine-tuning, become tightly coupled with specific training images. This follows the results of our theoretical understanding in Theorem 1. The attacker exploits this behavior to extract training data by performing the following steps:
  1. 1.
    Selection of Random Strings: The attacker identifies or generates a set of random strings $\mathcal{S}_{s}$ that were used during the victim’s fine-tuning process. These strings act as unique keys that the model associates with specific images.
  2. 2.
    Image Generation: For each random string $s\in\mathcal{S}_{s}$, the attacker generates $N_{G}$ images by querying the fine-tuned model:
    $\bm{x}_{i}^{(j)}\sim p_{\theta}(\bm{x}\mid s),\quad j=1,\dots,N_{G}$ (34)
    The repetition of image generation for the same string allows the attacker to assess the consistency of the outputs.
  3. 3.
    Consistency Analysis: The attacker computes the mean image $\overline{\bm{x}}_{s}$ and the sample variance $\sigma_{s}^{2}$ for the generated images associated with each string $s$:
    $\overline{\bm{x}}_{s}$ $=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i}^{(j)}$ (35)
    $\sigma_{s}^{2}$ $=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i}^{(j)}-\overline{\bm{x}}_{s}\|^{2}$ (36)
    A low variance $\sigma_{s}^{2}<\tau$ indicates that the generated images are highly similar, suggesting that the model has memorized the association between the string $s$ and a specific training image $\bm{x}_{s}$.
  4. 4.
    Data Extraction: If the variance $\sigma_{s}^{2}$ is below the threshold $\tau$, the attacker considers $\overline{\bm{x}}_{s}$ as an extracted training image and adds it to the dataset $\mathcal{D}_{\text{extracted}}$:
    $\mathcal{D}_{\text{extracted}}\leftarrow\mathcal{D}_{\text{extracted}}\cup\{\overline{\bm{x}}_{s}\}$ (37)
This method effectively leverages the deterministic mapping between random strings and training images induced by memorization, allowing the attacker to reconstruct the training dataset without direct access.
G.3 Algorithmic Implementation
The Extended SIDE method is formalized in Algorithm 2. The algorithm outlines the steps the attacker takes to generate images, compute the variance, and extract the training data based on the consistency of generated images. Since the attacker cannot compute similarity to the training data, they detect memorization by analyzing the diversity of generated images conditioned on the same random string $s$. The attacker identifies memorization by detecting low variance in the images generated from the same random string $s$. A low variance $\sigma_{s}^{2}$ indicates that the model consistently generates similar images for $s$, suggesting that it has memorized specific training data associated with $s$. The Extended SIDE is shown in Algorithm 2
G.4 Experiments
To verify the effectiveness of the Extended SIDE, we conduct experiments on the subset of the Laion-2B dataset, finetuning the Stable Diffusion v1.5 using the LORA[18] method. We use the following metrics to evaluate the attack results: mean SSCD (M-SSCD), AMS (mid-similarity) and LIPIPS. The results are shown in Tab. 4. We use the generated images of original prompts as the Baseline method. The extracted examples are shown in Figure 8. As the results show, our method could effectively extract training data from fine-tuned DPMs.
Table 4: Performance Comparison for Extended SIDE method
Dataset Method M-SSCD AMS LIPIPS
Subset of Laion-2B Baseline 0.028 0.000 0.892
Extended SIDE(ours) 0.467 0.672 0.809
As presented in Tab. 4, the Extended SIDE method outperforms the baseline across all evaluation metrics:
  • M-SSCD: The SIDE method achieves an M-SSCD of 0.467 compared to the baseline’s 0.028. This significant increase indicates that SIDE extracts images with higher structural similarity to the training data, demonstrating effective memorization.
  • AMS: SIDE records an AMS of 0.672, markedly higher than the baseline’s 0.000. This suggests that SIDE successfully captures a substantial number of moderately similar images, enhancing the diversity and relevance of the extracted data.
  • LIPIPS: SIDE exhibits a lower LIPIPS of 0.809 compared to the baseline’s 0.892. The difference is less pronounced. This is because the primary purpose of LIPIPS is to evaluate perceptual differences between images, rather than structural or training-specific similarity. It measures how humans perceive image similarity based on deep features from pre-trained networks, making it effective for detecting perceptual distortions or quality changes but not for identifying whether an image was directly memorized from the training data.
Visual examples Figure 8 showcases examples of extracted images using the SIDE method alongside the corresponding training images. The visual comparison underscores the high degree of similarity achieved by SIDE, affirming its efficacy in extracting memorized training data.
Figure 8: Examples of Extracted Images using SIDE Method. Top Row: Corresponding training images from the Laion-2B subset. Bottom Row: Images generated by querying the fine-tuned Stable Diffusion model with random strings using the SIDE method. The high visual similarity demonstrates the method’s effectiveness in extracting training data.
The Extended SIDE method effectively extracts training data from fine-tuned DPMs by exploiting the memorization induced by random strings during the fine-tuning process. The quantitative results demonstrate that SIDE significantly enhances the extracted images’ structural and mid-range similarity compared to the baseline. This indicates that SIDE can reliably reproduce specific training images by leveraging the unique associations between random strings and their corresponding images.