### v6

# appAppendix References

Abstract
As diffusion probabilistic models (DPMs) become the mainstream framework for Generative AI (GenAI), understanding their memorization behavior is critical for assessing risks of data leakage, copyright infringement, and ensuring trustworthy deployment. Previous studies indicate that conditional DPMs are more prone to memorization compared to unconditional DPMs, with the majority of extraction techniques focusing on conditional models. However, studying unconditional DPMs remains crucial, as they serve as the foundation for advanced models such as Stable Diffusion. Understanding their memorization behavior can provide valuable insights into the broader characteristics of DPMs. In this work, we propose Surrogate condItional Data Extraction (SIDE), a novel data extraction method for unconditional DPMs. SIDE employs a time-dependent classifier trained on generated data as surrogate conditions to extract training samples. Empirical results show that SIDE extracts training data effectively in challenging scenarios where existing methods fail, with an average performance gain of over 50% on the CelebA dataset across different scales. We also provide a unified theoretical framework explaining memorization in both conditional and unconditional DPMs and why SIDE works. Our findings contribute to a deeper understanding of memorization in GenAI and provide practical guidance for mitigating associated risks.
1 Introduction
Diffusion probabilistic models (DPMs) [21, 38, 42] are a family of powerful generative models that learn data distributions by progressively corrupting data through a forward diffusion process and then reconstructing it via a reverse diffusion process. Due to their exceptional capability to capture complex data distributions, DPMs have become foundation models for many pioneering Generative Artificial Intelligence (GenAI) products, including Stable Diffusion [33], DALL-E 3 [4], and Sora [5].
However, the growing adoption of DPMs raises concerns about data memorization — the risk of memorizing raw training samples. This may result in the generation of duplicated rather than novel content, increasing the risk of data leakage, privacy violations, and copyright infringement. [39, 40, 2, 13]. For instance, Stable Diffusion has been criticized as a “21st-century collage tool” that remixes copyrighted works from artists whose data was used for training [6]. Moreover, data memorization enables data extraction attacks, a type of privacy attack aimed at recovering training data from trained models. Recent work by [7] demonstrated the feasibility of extracting training data from DPMs like Stable Diffusion [33], highlighting significant privacy and copyright risks.
Existing studies have explored data memorization in DPMs, revealing that conditional DPMs exhibit a higher tendency to memorize training data than unconditional DPMs [40]. For example, [18] found that conditioning on random-labeled data significantly increases memorization, while unconditional models exhibit much lower memorization rates. Despite these findings, investigating data extraction attacks on unconditional DPMs remains crucial for two reasons: (1) unconditional DPMs form the foundation of conditional models like Stable Diffusion, which adapt pre-trained unconditional DPMs for conditional generation [33, 48]; and (2) understanding memorization in unconditional DPMs provides insights into the broader memorization mechanisms of DPMs.
In this work, we propose Surrogate condItional Data Extraction (SIDE), a novel method for extracting training data from unconditional DPMs. SIDE employs a time-dependent classifier trained on generated data as surrogate conditions to effectively extract training samples. We empirically validate SIDE on CIFAR-10, multiple scales of the CelebA and ImageNet, demonstrating its superiority over baseline methods by achieving over 50% higher effectiveness on average. Examples of extracted images are shown in Figure 1. Additionally, we introduce a divergence to measure memorization in DPMs and present a theoretical analysis explaining: (1) why conditional DPMs memorize more data (even labelled with random labels), (2) why SIDE is effective for extracting training data.
In summary, our main contributions are:
  • We propose SIDE, a novel method that leverages implicit labels provided by a time-dependent classifier to extract training data from unconditional DPMs.
  • We introduce a memorization divergence measure and provide a theoretical analysis that explains the impact of conditioning in DPMs and the effectiveness of SIDE.
  • We empirically validate the effectiveness of SIDE on CIFAR-10, CelebA and ImageNet, demonstrating it outperforms baseline methods by over 50% on average.
Figure 1: Examples of training images (top) and corresponding extracted images (bottom) from a DDPM trained on a subset of CelebA using our SIDE method.
2 Related Work
Diffusion Probabilistic Models
DPMs [38] have achieved state-of-the-art performance in image and video generation, as exemplified by models such as Stable Diffusion [33], DALL-E 3 [4], Sora [5], Runway [33] and Imagen [35]. These models excel on various benchmarks [15]. DPMs can be interpreted from two perspectives: 1) score matching [42], where model learns the gradient of data distribution [43], and (2) denoising diffusion [21], where Gaussian noise is added to clean images over multiple time steps, and the model is trained to reverse this process. For conditional sampling, [15] introduced classifier guidance to steer the denoising process, while [22] proposed classifier-free guidance, enabling conditional generation without explicit classifiers.
Memorization in Diffusion Models
Early research on memorization primarily focused on language models [8, 25], which later inspired subsequent studies on DPMs [40, 18, 31, 36, 16, 1, 16, 27, 47, 3, 17]. Notably, [39] found that 0.5-2% of generated images duplicate training samples, a result corroborated by [7] through more extensive experiments on both conditional and unconditional DPMs. Further studies [40, 18] linked memorization to model conditioning, showing that conditional DPMs are more prone to memorization. To address memorization, several methods have been proposed for detection and mitigation. For example, [46] introduced a method to detect memorization-triggering prompts by analyzing the magnitude of text-conditional predictions, achieving high accuracy with minimal computational overhead. [32] proposed metrics based on cross-attention patterns in DPMs to identify memorization. On the mitigation side, [9] developed anti-memorization guidance to reduce memorization during sampling, while [32] modified attention scores or masked summary tokens in the cross-attention layer. [46] minimized memorization by controlling prediction magnitudes during inference.
Despite these advances, the current research mostly focuses on conditional DPMs [7, 45], as extracting data from unconditional DPMs has been found to be extremely more challenging [18]. However, investigating unconditional DPMs remains critical because 1) they form the foundation of conditional models like Stable Diffusion, which adapt them for conditional generation, and 2) understanding memorization in unconditional DPMs provides insights into the broader behavior of DPMs. In response to these challenges, we propose SIDE, a novel data extraction method for unconditional DPMs, and provide a unified theoretical framework to explain memorization in both conditional and unconditional DPMs.
3 Surrogate Conditional Data Extraction
Threat Model
In this work, we focus on unconditional DPMs and assume that the attacker has access to the target DPM as well as a classifier trained on the same or similar dataset. The attacker’s objective is to extract the original training samples from the target DPM. In cases where the classifier is unavailable, the attacker can leverage generated data to train one, following a standard training procedure, potentially with certain necessary annotations. It is important to note that the classifier does not leak training data, as we will empirically demonstrate in Sec. 5.3. Its sole purpose, in this context, is to generate implicit labels, which can alternatively be achieved through clustering techniques.
3.1 Intuition of SIDE
Previous research [18, 40] have shown that conditional DPMs are more susceptible to memorization due to their reliance on explicit labels (e.g., class labels or text conditions). In contrast, unconditional DPMs lack access to these explicit labels. However, they still learn to form representation clusters from the training data, even though these clusters are not explicitly defined [10]. We term these representation clusters as implicit labels. By formalizing the clustering patterns within the training data, it becomes possible to construct implicit labels and condition unconditional DPMs on these labels to facilitate training data extraction.
In the following, we detail the construction of implicit labels for unconditional DPMs, their conversion into explicit labels, and the subsequent use of these explicit labels for training data extraction.
3.2 Constructing Implicit Labels
Intuitively, implicit labels can be generated during the sampling process of the diffusion model using a classifier. This classifier can be trained on the same/similar dataset as the diffusion model. When such a classifier is unavailable, representation clusters extracted by a pre-trained feature extractor (e.g., the CLIP image encoder) can serve as implicit labels. We assume that an implicit label $y_{I}$ is learned by the target unconditional DPM, with its sampling process described as:
$\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\Big{(}\nabla_{\bm{x}}\log p_{\theta}^{t}({\bm{x}})+$
$\nabla_{\bm{x}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\Big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w,$ (1)
where ${\bm{x}}$ represents the state vector, $f(x,t)$ denotes the drift coefficient, $g(t)$ is the diffusion coefficient, $\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$ denotes the gradient of model $p_{\theta}$ given ${\bm{x}}$ at time $t$, and $\mathrm{d}w$ corresponds to the increment of the Wiener process.
To approximate the gradient in the sampling process, we use the classifier that generates the implicit labels. However, neural network classifiers are prone to miscalibration [19], which can affect the sampling process. To address this, we introduce a hyperparameter $\lambda$ to calibrate the classifier’s probability output using a power prior:
$p^{t}_{\theta}\left({\bm{x}}|y_{I}\right)\propto p^{t\lambda}_{\theta}\left(y_{I}|{\bm{x}}\right)p^{t}_{\theta}\left({\bm{x}}\right).$ (2)
This leads to a classifier-conditioned sampling process:
$\mathrm{d}{\bm{x}}=$ $\Big{[}f({\bm{x}},t)-g(t)^{2}\big{(}\nabla_{{\bm{x}}}\log p_{\theta}^{t}({\bm{x}})$
$+\lambda\nabla_{{\bm{x}}}\log p_{\theta}^{t}(y_{I}|{\bm{x}})\big{)}\Big{]}\mathrm{d}t+g(t)\mathrm{d}w.$ (3)
This approach differs from [15], which introduced classifier guidance for improving sample quality under the assumption that with $Z$ being a constant. However, this assumption only holds when $\lambda=1$, as $Z$ is explicitly dependent on the ${\bm{x}}_{t}$ (the $t$-th step of sampling image ${\bm{x}}$ ) when $\lambda\neq 1$. Our derivation solves this issue by redefining $p^{t}_{\theta}\left(y|{\bm{x}}\right)$ using the power prior.
Figure 2: An illustration of our proposed time-dependent classifier training method on a pseudo-labeled synthetic dataset.
3.3 Training Time-dependent Classifier
In Eq. (3.2) above, the classifier is represented as $p_{\theta}^{t}(y|{\bm{x}})$, where $t$ denotes the timestep, indicating its time-dependent nature. However, we only have access to a time-dependent classifier. To address this limitation, we propose a method named Time-Dependent Knowledge Distillation (TDKD) to train a time-dependent classifier based on synthetic data generated at different sampling steps. The distillation process is illustrated in Figure 2. TDKD enables the classifier to provide time-dependent guidance during sampling through a two-step process:
Step 1: The network architecture is modified to incorporate time-dependent inputs. The structure of the time-dependent module and the specific modifications are detailed in Appendix B.
Step 2: A synthetic dataset with pseudo labels is generated to distill knowledge from the time-independent classifier to its time-dependent counterpart.
At Step 2, we can leverage the target DPM to generate a synthetic dataset, following established generative data augmentation techniques [12, 11]. Pseudo labels for the synthetic dataset are then generated using the time-independent classifier trained on the original dataset. Finally, the time-dependent classifier is trained on the labeled synthetic dataset. The training objective is defined as:
$\mathcal{L}_{distill}=D_{KL}\left(p_{\theta}(y_{I}|{\bm{x}}),p_{\theta}^{t}(y_{I}|{\bm{x}}_{t})\right)$ (4)
3.4 Overall Procedure of SIDE
With the trained time-dependent classifier $p_{\theta}^{t}(y|{\bm{x}}_{t})$ and the target DPM, our SIDE method extracts training data from the model through a conditional generation process, as described in Algorithm 1. Given the label $y=c$, the complete procedure of SIDE is as follows. It begins by selecting a set of values for $\lambda$, denoted as $\mathcal{S}_{\lambda}$ (Line 1). For each $\lambda$, the target DPM is required to generate $N_{G}$ data samples (Lines 18 - 27). Then, at each sampling timestep $t$, the gradient $\nabla_{{\bm{x}}_{t}}CE(c,p_{\theta}^{t}(y|{\bm{x}}_{t}))$ (Line 22) is computed, where $CE(\cdot)$ is the cross-entropy loss. This gradient, together with the target DPM, is then used to guide the reverse diffusion process. For each generated image, a similarity score is computed against every ground-truth training images to identify the most closely matching original image.
Using the similarity scores obtained for the extracted images, a performance metric can then be calculated to evaluate the effectiveness of the extraction process. Existing methods predominantly rely on $L_{p}$ distance metrics to identify pixel-wise similar samples. However, such metrics are insufficient for capturing semantically meaningful similarities. In the experiment section, we will introduce a set of performance metrics to facilitate a more comprehensive evaluation of the extraction methods.
Algorithm 1 SIDE
1:Input: DPM $p_{\theta}(\bm{x})$, classifier $p_{\theta}(y_{I}|\bm{x})$, $\mathcal{S}_{\lambda}$, $N_{G}$, label $c$, $N_{syn}$, $T$, Denoise Step $DS(\bm{x}_{t},t,\bm{s})$
2:Output: Extracted data $\mathcal{D}_{ext}$, classifier $p_{\theta}^{t}(y_{I}|\bm{x}_{t})$
3:Classifier Training:
4:Step 1: Architecture
5:Modify $p_{\theta}(y|\bm{x})$ to $p_{\theta}^{t}(y|\bm{x}_{t})$ for time input
6:Step 2: Distillation
7:for $i=1$ to $N_{syn}$ do $\triangleright$ Generate synthetic data
8:    $\bm{x}_{T}\sim\mathcal{N}(0,I)$
9:    for $t=T$ down to $1$ do
10:         $\bm{x}_{t-1}\leftarrow DS(\bm{x}_{t},t,s_{\theta}(\bm{x}_{t},t))$
11:    end for
12:    $y_{I}^{(i)}\leftarrow p_{\theta}(y_{I}|\bm{x}_{0}^{(i)})$triangleright Generate pseudo-labels
13:end for
14:Step 3: Train $p_{\theta}^{t}(y_{I}|\bm{x}_{t})$:
15:$\mathcal{L}_{distill}=\frac{1}{N_{syn}}\sum_{i=1}^{N_{syn}}D_{KL}(p_{\theta}(y_{I}|\bm{x}_{0}^{(i)}),p_{\theta}^{t}(y_{I}|\bm{x}_{t}^{(i)}))$
16:Data Extraction:
17:Initialize $\mathcal{D}_{ext}\leftarrow\emptyset$
18:for $\lambda\in\mathcal{S}_{\lambda}$ do
19:    for $i=1$ to $N_{G}$ do
20:         $\bm{x}_{T}\sim\mathcal{N}(0,I)$
21:         for $t=T$ down to $1$ do
22:             $g_{t}\leftarrow\nabla_{\bm{x}_{t}}CE(c,p_{\theta}^{t}(y_{I}|\bm{x}_{t}))$
23:             $s_{\text{guided}}\leftarrow\nabla_{\bm{x}_{t}}\log p_{\theta}^{t}(\bm{x}_{t})-\lambda g_{t}$
24:             $\bm{x}_{t-1}\leftarrow DS(\bm{x}_{t},t,s_{\text{guided}})$
25:         end for
26:         Append $\bm{x}_{0}$ to $\mathcal{D}_{ext}$
27:    end for
28:end for
29:return $\mathcal{D}_{ext}$, $p_{\theta}^{t}(y_{I}|\bm{x}_{t})$
4 Theoretical Analysis
In this section, we first introduce a Kullback-Leibler (KL) divergence-based distribution measure to quantify the degree of memorization in generative models. Building on this measure, we provide a theoretical explanation for data memorization in conditional diffusion probabilistic models (DPMs) and elucidate why SIDE is effective for unconditional DPMs.
4.1 Distributional Memorization Measure
There are various approaches to measure the memorization effect of a generative model. One approach involves comparing each generated sample with the raw training samples in a sample-by-sample manner, as previously mentioned. While such measures (e.g., $L_{p}$ distances) are effective for evaluating data extraction performance, they are insufficient for analyzing the overall memorization effect of the model. To characterize the model-level memorization effect with respect to the training data distribution and to support our theoretical analysis, we introduce the following distributional memorization measure.
To assess the model’s memorization ability with respect to the overall data distribution while also accounting for the memorization of each individual data point, it is natural to consider the Kullback-Leibler (KL) divergence between the uniform empirical distribution over the dataset $\mathcal{D}$, defined as $\frac{1}{|\mathcal{D}|}\sum_{x_{i}\in\mathcal{D}}\delta(x_{i})$ with $\delta(\cdot)$ is the Dirac delta function, and the distribution $p$ induced by the model’s generated samples. Here, the $\delta(\cdot)$ function enforces a point-wise memorization measure, which quantifies how closely the generated samples align with each individual data point in the original distribution. A smaller KL divergence indicates that the model exhibits stronger memorization of the training data points. However, for a continuous probability distribution $p$, computing such a KL divergence directly is infeasible. To address this, we approximate each Dirac delta function by a normal distribution with a small variance in the following definition.
Definition 1 (Memorization Divergence).
Given a generative model $p_{\theta}$ with parameters $\theta$ and training dataset $\mathcal{D}=\{{\bm{x}}_{i}\}_{i=1}^{N}$, the degree of divergence between $p_{\theta}$ and distribution of training dataset is defined as:
$\mathcal{M}(\mathcal{D};p_{\theta},\epsilon)$ $=D_{\mathrm{KL}}(q_{\epsilon}\|p_{\theta})$ (5)
with $\text{with }q_{\epsilon}(x)=$ $\frac{1}{N}\sum_{x_{i}\in\mathcal{D}}\mathcal{N}(x|x_{i},\epsilon^{2}I),$
where ${\bm{x}}_{i}\in\mathbb{R}^{d}$ denotes the $i$-th training sample, $N$ is the total number of training samples, $p_{\theta}({\bm{x}})$ represents the probability density function (PDF) of the generated samples, and $\mathcal{N}(x|x_{i},\epsilon^{2}I)$ is the normal distribution with mean ${\bm{x}}_{i}$ and covariance matrix $\epsilon^{2}I$.
Note that in Eq. (5), a smaller value of ${\mathcal{M}}({\mathcal{D}};p_{\theta},\epsilon)$ indicates greater overlap between the two distributions, reflecting stronger memorization. As $\epsilon$ approaches 00, the measured memorization divergence becomes more accurate. In fact, the normal distribution $\mathcal{N}(x|x_{i},\epsilon^{2}I)$ can be replaced by any family of continuous distributions $\hat{q}(x|x_{j},\epsilon)$ that satisfies two conditions: 1) symmetric with respect to $x$ and $x_{j}$(i.e., $\hat{q}(x|x_{j},\epsilon)=\hat{q}(x_{j}|x,\epsilon)$) and 2) convergence to $\delta(x_{i})$ in distribution. This substitution does not affect the result of Theorem 1 in Sec. 4.2.
While concerns may arise regarding the influence of $\epsilon$ on the divergence, this divergence is primarily intended for comparative analysis when $\epsilon$ is sufficiently small. When comparing the memorization divergence across different models, the $\epsilon$ will not influence relative differences, as demonstrated in Appendix A.
4.2 Theoretical Analysis
Building on the memorization divergence measure, here we provide a theoretical analysis to explain why conditional DPMs exhibit stronger memorization effect. Our analysis centers on the concept of informative labels, which represent information capable of partitioning a dataset into multiple disjoint subsets. We show that DPMs, conditioned on informative labels, exhibit stronger memorization effects.
Informative Labels
The concept of informative labels has been previously discussed as class labels [18]. In this work, we extend this notion by introducing a more general definition that encompasses includes both class labels and random labels as special cases. Formally, we define an informative label as follows.
Definition 2 (Informative Label).
Let ${\mathcal{Y}}$ be an attribute of the data that takes values in $\{y_{i}\}_{i=1}^{C}$. We define ${\mathcal{Y}}$ as an informative label if it enables the partitioning of the dataset into several mutually disjoint subsets $\{{\mathcal{D}}_{i}\}_{i=1}^{C}$, where each subset corresponds to a distinct value of ${\mathcal{Y}}$.
In the above definition, informative labels are not restricted to conventional class labels; they can also include text captions, features, or cluster information that can be used to group training samples into subsets. The definition emphasizes that an informative label must be capable of distinguishing a subset of samples from others. An extreme case occurs when all samples share the same label, rendering the label non-informative. According to our definition, both class-wise and random labels are special cases of informative labels. Informative labels can be either explicit, such as class labels, random labels, or text captions, or implicit, such as salient features or clusters.
Next, we present our main theoretical result on the memorization mechanism of conditional DPMs and explain why SIDE is effective. Let ${\mathcal{D}}_{i}$ denote the subset consisting of data with the informative label ${\mathcal{Y}}=y_{i}$. Additionally, we denote the data distribution in the original dataset ${\mathcal{D}}$ by $p$. Similarly, for each attribute $y_{i}$, we use $p_{i}$ to denote the data distribution corresponding to that attribute.
Theorem 1.
If a generative model $p_{\theta_{i}}$ matches the target distribution $p_{i}$ almost everywhere for the informative label $y_{i}$, i.e., $TV(p_{i},p_{\theta_{i}})=0$, then with probability 1:
$\lim_{\epsilon\to 0}\lim_{|{\mathcal{D}}_{i}|\to\infty}\left({\mathcal{M}}({\mathcal{D}}_{i};p_{\theta_{i}},\epsilon)-{\mathcal{M}}({\mathcal{D}}_{i};p_{\theta},\epsilon)\right)\leq 0,$ (6)
where $TV(\cdot)$ is the total variance distance, $p_{\theta_{i}}$ and $p_{\theta}$ denote the distribution of generated data for model trained on data labeled $y_{i}$ and the whole dataset. The equality holds if and only if $TV(p,p_{i})=0$.
The proof for Theorem 1 is in Appendix A. The theorem implies that any form of informative labels, including class labels and random labels, can enhance the memorization effect of conditional generative models. Conditional models exhibit stronger memorization by isolating data subsets through explicit informative labels, allowing focused training on each group’s unique patterns, which sharpens their ability to memorize specific training examples. SIDE succeeds by uncovering and exploiting the unconditional DPM’s inherent but implicit informative labels, using classifier guidance to concentrate generation on specific memorized features.
Table 1: The AMS (%) and UMS (%) results at low, mid, and high similarity levels. ‘Random’ is the baseline method proposed in [7], while OL-TI is SIDE without TDKD training. We also compare SIDE using classifiers trained on the original (SIDE-O) vs. disjoint auxiliary (SIDE-A) datasets. The similar performance between SIDE-O and SIDE-A suggests there is no training data leaking via the classifier.
Dataset Method Low Similarity Mid Similarity High Similarity
AMS(%) UMS(%) AMS(%) UMS(%) AMS(%) UMS(%)
CIFAR-10 Random 2.470 1.770 0.910 0.710 0.510 0.420
OL-TI 2.460 1.780 0.800 0.680 0.420 0.370
SIDE-A 5.250 2.020 2.300 0.880 1.620 0.640
SIDE-O 5.320 2.050 2.490 0.860 1.770 0.560
CelebA-HQ-FI Random 11.656 2.120 0.596 0.328 0.044 0.040
OL-TI 2.649 0.744 0.075 0.057 0.005 0.005
SIDE-A 15.010 2.624 1.310 0.554 0.090 0.082
SIDE-O 15.172 2.342 1.115 0.444 0.054 0.044
CelebA-25000 Random 5.000 4.240 0.100 0.100 0.000 0.000
OL-TI 0.164 0.152 0.000 0.000 0.000 0.000
SIDE-A 8.712 6.802 0.234 0.234 0.010 0.010
SIDE-O 8.756 6.940 0.224 0.212 0.012 0.012
CelebA Random 1.953 1.895 0.000 0.000 0.000 0.000
OL-TI 1.145 1.125 0.000 0.000 0.000 0.000
SIDE-A 4.682 4.706 0.098 0.098 0.000 0.000
SIDE-O 4.883 4.766 0.078 0.078 0.000 0.000
ImageNet Random 0.000 0.000 0.000 0.000 0.000 0.000
OL-TI 0.000 0.000 0.000 0.000 0.000 0.000
SIDE-A 0.152 0.152 0.076 0.076 0.000 0.000
SIDE-O 0.114 0.114 0.038 0.038 0.000 0.000
Figure 3: A comparison between the original training images (top row) and extracted images (bottom row) by our SIDE method. The matches are classified into three categories: low similarity (SSCD score $<$ 0.5), mid similarity (SSCD score between 0.5 and 0.6), and high similarity (SSCD score $>$ 0.6). This highlights varying degrees of semantic resemblance among the image pairs.
5 Experiments
In this section, we first introduce the experimental setup and performance metrics, followed by the main evaluation results. We also show that the memorized images are not derived from the classifier. Additionally, we conduct an ablation study and hyperparameter analysis to provide deeper insights into the working mechanism of SIDE.
5.1 Experimental Setup
We evaluate the effectiveness of our method on five datasets: 1) CelebA-HQ-Face-Identity (CelebA-HQ-FI) [30], which consists of 5478 images with a resolution of 128$\times$128; 2) a subset of the CelebA (CelebA-25000) [28], containing 25,000 images with a resolution of 128$\times$128; 3) CIFAR-10, containing 50,000 images with a resolution of 32$\times$32; 4) CelebA [28], which contains 22059 images with a resolution of 128$\times$128; and 5) ImageNet [14], which contains $\approx$ 14 million images with a resolution of 256$\times$256. All the images are normalized to the range of [-1,1]. We use the AdamW optimizer [29] with a learning rate of 1e-4 to train the time-dependent classifier. We train denoising DPMs with a discrete denoising scheduler (DDIM [41]) using the HuggingFace implementation [44]. All DPMs are trained with batch size 64. We train the models for 258k steps ($\approx$ 3000 epochs) on CelebA-HQ-FI, 390k steps ($\approx$ 1000 epochs) on CelebA-25000 and CelebA, 1600k steps ($\approx$ 2048 epochs) on CIFAR-10, and 1980K steps on ImageNet, respectively. We use ResNet34 [20] as the pseudo-labeler. For ImageNet, the DPM is trained on the entire dataset. However, due to the scale of ImageNet, we only evaluate on a smaller subset known as ImageNette [23].
5.2 Image-level Performance Metrics
As previously mentioned, determining whether an extracted image is a memorized copy of a training image is challenging, and $L_{p}$ distances in pixel space are ineffective when the image is semantically similar but not an exact copy. Prior research [39, 18] addresses this by using the 95th percentile Self-Supervised Descriptor for Image Copy Detection (SSCD) score for image copy detection. However, the 95th percentile SSCD score has three limitations: 1) it does not measure the uniqueness of memorized images; 2) it may underestimate the number of memorized samples when cut at the 95th percentile; and 3) it does not account for different types of memorization. To address these issues, we propose two new memorization scores: 1) Average Memorization Score (AMS) and 2) Unique Memorization Score (UMS). The AMS and UMS scores are defined as follows:
$AMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{\sum_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}{\mathcal{F}\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)}}{N_{G}}$ (7)
$UMS\left({{\mathcal{D}}}_{1},{\mathcal{D}}_{2},\alpha,\beta\right)=\frac{|\bigcup_{{\bm{x}}_{i}\in{\mathcal{D}}_{1}}\phi\left({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta\right)|}{N_{G}},$ (8)
where ${{\mathcal{D}}}_{1}$ is the generated dataset, ${\mathcal{D}}_{2}$ is the training dataset, and $\alpha$, $\beta$ are thresholds for image similarity scoring. $\mathcal{F}({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ serves as a binary check for whether any training sample meets the similarity/distance criteria. $\phi({\bm{x}}_{i},{\mathcal{D}}_{2},\alpha,\beta)$ provides the specific indices of those training samples that meet the similarity/distance criteria. Mathematically, they can be formulated as:
$\mathcal{F}(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\mathds{1}\Big{[}\max_{x_{j}\in{\mathcal{D}}_{2}}\gamma(x_{i},x_{j})\in[\alpha,\beta]\Big{]}$ (9)
$\phi(x_{i},{\mathcal{D}}_{2},\alpha,\beta)=\{j:x_{j}\in{\mathcal{D}}_{2},\gamma(x_{i},x_{j})\in[\alpha,\beta]\}$ (10)
where $\gamma$ denotes the similarity/distance function. For low-resolution datasets, we use the normalized $L_{2}$ distance as $\gamma$, following [7], while for high-resolution datasets, we use the SSCD score as $\gamma$.
The AMS averages the similarity scores across extracted images, ensuring that memorized images are not overlooked. In contrast, the UMS quantifies distinct memorized images by evaluating unique matches, thereby accounting for the uniqueness of the memorized images. By further categorizing the two scores into three levels—low, mid, and high—we obtain more comprehensive measurements for different types of memorization.
In our experiments, the thresholds for SSCD are set to $\alpha=0.4$ and $\beta=0.5$ for low similarity, $\alpha=0.5$ and $\beta=0.6$ for mid similarity, and $\alpha=0.6$ and $\beta=1.0$ for high similarity. (examples of different thresholds are shown in Figure 3) The thresholds for the normalized $L_{2}$ distance [7] are set to $\alpha=1.5$ and $\beta=10$ for low similarity, $\alpha=1.4$ and $\beta=1.5$ for mid similarity, and $\alpha=1.35$ and $\beta=1.4$ for high similarity.
Relation to Existing Metrics  While [7, 9] introduced metrics similar to AMS and UMS, they did not account for varying levels of similarity, which is crucial for assessing different types of copyright infringement, such as character or style copying [26, 34, 37]. Additionally, the UMS considers the number of extracted images $N_{G}$, a factor overlooked in [7]. The significance of $N_{G}$ lies in its non-linear impact on the UMS. Let $p_{\gamma}(i)$ be the probability that image $i$ is memorized, and $M$ be the size of the dataset. The expected number of uniquely memorized images is given by: $\mathbb{E}[N_{\mathrm{umem}}]=\sum_{i=1}^{M}{1-\left(1-p_{\gamma}\left(i\right)\right)^{N_{G}}}$, revealing that varying $N_{G}$ affects UMS non-linearly. This means that we should compare UMS under the same $N_{G}$. Note that the AMS and UMS scores are also different from the memorization divergence measure defined in Eq. (1), which is a distributional measure rather than an individual measure.
Table 2: The coefficient and $R^{2}$ results of a fitted linear model between the number of classes vs. AMS/UMS. $R^{2}$ measures how well a regression model fits the data (1 is perfect fitting).
Relationship Coefficient $R^{2}$
Class vs. AMS 7.4 ×$\times 10^{-5}$ 0.637
Class vs. UMS 6.1 ×$\times 10^{-5}$ 0.483
5.3 Main Results
We compare SIDE with the Random baseline proposed by Carlini et al.[7] and a variant of SIDE that substitutes the time-dependent classifier with a standard (time-independent) classifier. Here, ”TD” refers to the time-dependent classifier trained using our proposed TDKD method, ”TI” denotes the time-independent classifier, and ”OL” indicates training with original dataset labels.
The Random baseline generates images directly using the target unconditional DPM, as described in [7]. We average the results across various values of $\lambda$ (defined in Eq. (3.2)), ranging from 5 to 10, with a detailed analysis provided in Section 5.3. It is important to note that $\lambda=0$ corresponds to the Random baseline.
To evaluate SIDE, for each $\lambda$, including $\lambda=0$, we generate 51,200 images for CelebA-HQ-FI, 50,000 images for CelebA-25000, 10,000 images for CIFAR-10, 5,120 images for CelebA and 2,560 images for ImageNet. The results are presented in Table 4.
Figure 4: Hyper-parameter ($\lambda$) analysis on CelebA-HQ-FI. For high similarity, the best $\lambda$ for AMS and UMS are 16 and 13. For other similarity levels, the best $\lambda$ for AMS and UMS is 13.
Figure 5: The relationship between the number of classes and AMS/UMS. The fitted regression lines demonstrate a positive correlation in both metrics.
Effectiveness of SIDE
Table 4 summarizes the AMS and UMS results for the five datasets (results of 95th percentile SSCD are in Appendix C). SIDE shows strong performance in extracting training data across various datasets and similarity levels. Data extraction of unconditional DPM on large-scale datasets was considered impossible. But SIDE successfully extracts from these datasets.
On CIFAR-10, SIDE outperforms baselines at all similarity levels. For low similarity, it achieves 5.320% AMS (vs. Random’s 2.470%) and 2.05% UMS (vs. 1.780%). For mid similarity, SIDE reaches 2.490% AMS and 0.860% UMS, surpassing Random’s 0.910% AMS and 0.710% UMS. At high similarity, SIDE achieves 1.770% AMS and 0.560% UMS, outperforming Random’s 0.510% AMS and 0.420% UMS.
On CelebA-HQ-FI, SIDE boosts mid-level AMS by 87% to 1.115% and UMS by 37% to 0.444%, with a 20% average improvement across other levels. SIDE could extract 900 high, 11,500 mid, and 151,720 low similarity images from 1 million samples. On CelebA-25000, SIDE improves AMS and UMS by 75% and 63% for low similarity, and 124% and 112% for mid similarity, excelling at high-similarity memorized data extraction. For CelebA and ImageNet, SIDE successfully extracts training data where baselines fail.
Effectiveness of TDKD
As shown in Table 4, time-independent classifiers perform significantly worse than their time-dependent counterparts on the two high-resolution CelebA datasets, achieving only about 10% of the effectiveness of classifiers trained using our TDKD method. This discrepancy arises because time-independent classifiers produce accurate gradients only at the final timestep, whereas time-dependent classifiers deliver accurate gradients for each timestep. However, the performance gap narrows on the CIFAR-10 dataset due to its simplicity, consisting of only 10 classes. As a result, the gradients produced by the time-independent classifier are less susceptible to inaccuracies, thereby mitigating its inherent limitations.
No Training Data Leakage From the Classifier
A potential concern is that the extracted images might reflect artifacts introduced by the classifier rather than the target DPM itself. However, we argue that even if a time-dependent classifier were to expose its training images, those images would inherently originate from the training set of the target DPM, as the classifier is distilled on synthetic data generated by the DPM. Furthermore, as shown in Table 4, using a simple classifier reduces memorization compared to the baseline extraction method, indicating that the extracted memorized images do not primarily originate from the classifier. To further confirm our conjecture, we compare SIDE using classifiers trained on two different datasets: SIDE-A (an auxiliary dataset, disjoint from the DPM’s training data) and SIDE-O (the original training dataset). As shown in Table 4, SIDE-A and SIDE-O achieve similar results. This verifies that the classifier does not leak training data information, as SIDE only leverages it to label synthetic data rather than generate data samples.
Figure 6: Validation of $N_{G}$ ’s significance
Impact of the Classifier on SIDE
The classifier used to train SIDE is associated with a certain number of classes. Here, we explore the relationship between the number of classes (of the classifier) and the extraction performance at a low similarity level, using 1,200 images per class with $\lambda=5$ on the CelebA-HQ-FI dataset. As shown in Table 2 and Figure 5, the number of classes exhibits a strong positive correlation with both AMS and UMS. Increasing the number of classes enhances both metrics, with a more pronounced effect on AMS. The UMS values here differ significantly from those in Table 4 due to the smaller sample size per class (1,200 images) in this experiment. The stronger improvement in AMS arises because the classifier captures fine-grained details more effectively with more classes, leading to more accurate matches at lower similarity levels. In contrast, UMS is less affected, as it depends more on overall image diversity, which is constrained by the smaller sample size per class in this setting.
Importance of $N_{G}$ in UMS
The number of uniquely memorized samples in a dataset of size $M$ can be formulated as $\sum_{i=1}^{M}{1-\left(1-k_{i}^{N_{G}}\right)}$, where $k_{i}$ denotes the probability the $i$-th sample is extracted per trial. To empirically verify the importance of $N_{G}$ , we generate 1 million samples using a DPM trained on CelebA-HQ-FI. As shown in Figure 6, the theoretical and empirical results align closely, confirming that $N_{G}$ non-linearly influences UMS.
Hyper-parameter Analysis
Here, we test the sensitivity of SIDE to its hyper-parameter $\lambda$. We generate 50,000 images for each integer value of $\lambda$ within the range of [0, 50]. As shown in Figure 4, the memorization score increases at first, reaching its highest, then decreases as $\lambda$ increases. This can be understood from sampling SDE Eq. (3.2). Starting from 0, the diffusion models are unconditional. As $\lambda$ increases, the diffusion models become conditional, and according to Theorem 1, the memorization effect will be triggered. However, when $\lambda$ becomes excessively large, the generated images will overfit the classifier’s decision boundaries, resulting in reduced diversity and a failure to accurately reflect the underlying data distribution. Consequently, the memorization score decreases.
Limitation
While showing promising results on unconditional DPMs, our SIDE still relies on a reliable pseudo-labeler, such as a classifier trained on the auxiliary data or an accurate clustering method. We plan to reduce this dependency by exploring pre-trained image encoders in future work. Additionally, SIDE requires access to the original training data for performance evaluation, which may not be feasible for commercial models. Future research could focus on developing evaluation methods that do not depend on the original training data.
6 Conclusion
In this paper, we proposed a novel method SIDE for extracting training data from unconditional diffusion probabilistic models (DPMs) by constructing surrogate conditions. At its core, SIDE trains a time-dependent classifier to create these surrogate conditions. We empirically validated the effectiveness of SIDE on datasets of varying scales, including CelebA, ImageNet, and CIFAR-10. Additionally, we provide a theoretical analysis to explain the role of information labels in facilitating memorization in DPMs. Our work highlights the potential for memorization in unconditional DPMs and demonstrates that such memorized images can be extracted by constructing surrogate conditions. Our study advances the understanding of memorization mechanisms in DPMs and motivates the development of more effective memorization mitigation strategies, such as regularizing deep representations to be less informative.
Supplementary Material
Broader Impacts
Our work not only quantifies memorization effects in diffusion models, but also extends to deep generative models. This contribution is significant in enhancing our understanding of when and how models memorize training data, which is critical for addressing concerns about data privacy and model robustness. By providing a theoretical framework for conditional memorization, we pave the way for developing effective memorization mitigation strategies tailored for diffusion models. These advancements can lead to the design of more secure and trustworthy AI systems, reducing the risk of potential data leakage while fostering greater accountability in the use of generative technologies. Ultimately, our findings aim to empower researchers and practitioners to create models that better respect privacy, thus benefiting the wider AI community.
Appendix A Proof of Theorem1
By the assumption $TV(p_{i},p_{\theta_{i}})=0$ , we can replace $p_{\theta_{i}}$ with $p_{i}$ and $p_{\theta}$ with $p$ to simplify the notation without compromising correctness. Hence, we have:
$\mathcal{M}(\mathcal{D}_{i};p_{\theta_{i}},\epsilon)-\mathcal{M}(\mathcal{D}_{i};p_{\theta},\epsilon)=\frac{1}{N}\sum_{x_{k}\in\mathcal{D}_{i}}\int\mathcal{N}(x|x_{k},\epsilon^{2}I)\log{\frac{p(x)}{p_{i}(x)}}dx$ (11)
By the Strong Law of Large Numbers, as $N\to\infty$, with probability 1:
$\frac{1}{N}\sum_{x_{k}\in\mathcal{D}_{i}}\mathcal{N}(x|x_{k},\epsilon^{2}I)$ $\to\mathbb{E}_{y\sim p_{i}}\left[\frac{1}{(2\pi\epsilon^{2})^{d/2}}\exp{(-\frac{||y-x||^{2}}{2\epsilon^{2}})}\right]$ (12)
[…]
Now, we show term $L\to 0$ as $\epsilon\to 0$. By the continuity of $p_{i}$, for any $\eta>0$, select $r>0$ such that $|p_{i}(y)-p_{i}(x)|<\eta,\ \ \forall\ ||y-x||<r$. Then we can decomposite $L$ into two parts:
$|L|$ $\leq\frac{1}{(2\pi\epsilon^{2})^{d/2}}\left(\int_{||y-x||<r}+\int_{||y-x||\geq r}\right)|p_{i}(y)-p_{i}(x)|\exp{(-\frac{||y-x||^{2}}{2\epsilon^{2}})}dy$ (15)
[…]
$\to\eta\text{ \ \ as }\epsilon\to 0$ (17)
Due to the arbitrariness of $\eta$, $L\to 0$ as $\epsilon\to 0$, we obtain
$\lim_{\epsilon\to 0}\lim_{|{\mathcal{D}}_{i}|\to\infty}({\mathcal{M}}({\mathcal{D}}_{i};p_{\theta_{i}},\epsilon)-{\mathcal{M}}({\mathcal{D}}_{i};p_{\theta},\epsilon))=\int p_{i}(x)\log{\frac{p(x)}{p_{i}(x)}}dx=-D_{\mathrm{KL}}(p_{i}||p)\leq 0$ (18)
Appendix B Refinement ResNet block (Figure 7)
Figure 7: Refinement ResNet block with time-dependent module integration. This block diagram depicts the insertion of a time module within a conventional ResNet block architecture, allowing the network to respond to the data’s timesteps. Image $x_{BN}$ is the image processed after the first Batch Normalization Layer.
This section elaborates on our time module’s design principles and architectural rationale, which strategically integrates temporal dynamics into normalized feature spaces through a post-batch normalization framework.The integration of the time module directly after batch normalization within the network architecture is a reasonable design choice rooted in the functionality of batch normalization itself. Batch normalization standardizes the inputs to the network layer, stabilizing the learning process by reducing internal covariate shifts. The model can introduce time-dependent adaptations to the already stabilized features by positioning the time module immediately after this normalisation process. This placement ensures that the temporal adjustments are applied to a normalized feature space, thereby enhancing the model’s ability to learn temporal dynamics effectively.
Moreover, the inclusion of the time module at a singular point within the network strikes a balance between model complexity and temporal adaptability. This singular addition avoids the potential redundancy and computational overhead that might arise from multiple time modules. It allows the network to maintain a streamlined architecture while still gaining the necessary capacity to handle time-varying inputs.
Appendix C SSCD-related metrics
To comprehensively evaluate the proposed method, it’s crucial to present a discussion of the 95th percentile SSCD metric alongside our newly introduced metrics, AMS and UMS. Although the 95th percentile SSCD metric has its limitations, it still serves as a useful reference point for assessing the relative similarity between extracted images and the original training dataset. As shown in Table 3, we provide a comprehensive comparison of performance across different datasets.
Table 3: Evaluation of Data Extraction Performance using SSCD Percentiles. The table shows SSCD scores (Top 0.05% to Top 5.0%) for the Random baseline and SIDE methods (SIDE-A, SIDE-O) across CelebA-HQ-FI, CelebA-25000, CelebA, and ImageNet datasets.
Dataset Method Top 0.05% Top 0.1% Top 0.5% Top 1.0% Top 5.0%
CelebA-HQ-FI Random 0.604 0.584 0.524 0.498 0.433
SIDE-A 0.617 0.611 0.576 0.555 0.485
SIDE-O 0.639 0.618 0.567 0.544 0.485
CelebA-25000 Random 0.518 0.505 0.471 0.455 0.404
SIDE-A 0.534 0.525 0.500 0.485 0.439
SIDE-O 0.554 0.539 0.506 0.491 0.450
CelebA Random 0.487 0.477 0.452 0.439 0.404
SIDE-A 0.523 0.517 0.495 0.479 0.436
SIDE-O 0.513 0.508 0.471 0.456 0.419
ImageNet Random 0.330 0.316 0.287 0.275 0.250
SIDE-A 0.512 0.493 0.378 0.342 0.283
SIDE-O 0.462 0.419 0.362 0.342 0.291
Appendix D Cluster-Based SIDE (SIDE-C)
In the main text, we presented the SIDE algorithm using a classifier as a surrogate condition for simplicity and clarity in the core algorithm description. However, we also investigated the use of cluster-based surrogate conditions. This appendix details the formulation of the cluster-based SIDE approach and presents a performance analysis.
In our cluster-based approach, we first extract feature dataset from the synthetic dataset generated by the target DPM using SSCD encoder. This extracted data serves as the basis for our clustering. We then apply k-means clustering to this SSCD-extracted dataset, partitioning the data into 100 distinct clusters.These cluster assignments are then treated as implicit labels. Instead of using a classifier’s predicted class probabilities to guide the diffusion process, we use the cluster labels to train the time-dependent classifier.
In essence, we replace the class-conditional term in the gradient guidance with a cluster-conditional guidance term. This allows the diffusion model to be guided towards generating data points that are representative of specific clusters discovered within the SSCD-extracted data, effectively using these clusters as implicit, self-discovered categories for data extraction.
Table 4: Performance analysis of SIDE with cluster-based surrogate conditions. This table compares AMS (%) and UMS (%) in percentage for different methods, including baselines (Random, OL-TI) and SIDE variations (classifier-based: SIDE-O, SIDE-A; cluster-based: SIDE-C), across low, mid, and high similarity levels on CIFAR-10 and CelebA-25000 datasets.
Dataset Method Low Similarity Mid Similarity High Similarity
AMS(%) UMS(%) AMS(%) UMS(%) AMS(%) UMS(%)
CIFAR-10 Random 2.470 1.770 0.910 0.710 0.510 0.420
OL-TI 2.460 1.780 0.800 0.680 0.420 0.370
SIDE-A 5.250 2.020 2.300 0.880 1.620 0.640
SIDE-O 5.320 2.050 2.490 0.860 1.770 0.560
SIDE-C 7.830 2.730 3.830 1.190 2.610 0.760
CelebA-25000 Random 5.000 4.240 0.100 0.100 0.000 0.000
OL-TI 0.164 0.152 0.000 0.000 0.000 0.000
SIDE-A 8.712 6.802 0.234 0.234 0.010 0.010
SIDE-O 8.756 6.940 0.224 0.212 0.012 0.012
SIDE-C 20.500 11.400 1.842 1.164 0.000 0.000
SIDE-C (cluster-based SIDE) generally outperforms other methods, especially classifier-based SIDE variants (SIDE-O, SIDE-A), in data extraction, achieving significantly higher AMS and UMS on both CIFAR-10 and CelebA-25000, particularly at low and mid similarity levels. The cluster-based approach appears to offer a more effective guidance mechanism, potentially due to implicit labels derived from data-driven clusters. However, a notable exception is observed at high similarity for CelebA-25000, where SIDE-C’s performance dramatically drops to 0%, suggesting a potential failure mode, possibly related to cluster overfitting or interaction with strong similarity constraints, unlike the more consistent, though less performant, classifier-based SIDE. This highlights SIDE-C’s promise as a powerful data extraction technique, but also points to areas needing further investigation, particularly its robustness under high similarity conditions.
Appendix E Extended SIDE for Text-to-Image Diffusion Models
We extend the SIDE method for text-to-image diffusion models in a fine-tuning scenario. Instead of using a classifier as a surrogate condition to extract training data, we use random strings as surrogate conditions. The attack setting is simple and practical: the attacker can access the fine-tuned model and modify the text inputs in the text-to-image pairs. Notably, the attacker does not have access to the training data. We name this method as Extended SIDE.
E.1 Attack Setting
The attack setting for the Extended SIDE method involves two primary components: the victim’s fine-tuning process and the attacker’s access to the fine-tuned model. In the following, we provide a detailed description of each component.
  • Victim Fine-Tuning Process: The victim fine-tunes a text-to-image diffusion model, denoted as $p_{\theta}(\bm{x}\mid s)$, using a dataset of image-text pairs $\mathcal{D}_{train}=\{(s_{i},\bm{x}_{i})\}_{i=1}^{N}$. In this context $s_{i}\in\mathcal{S}_{s}$ are random strings crafted by the attacker to serve as surrogate text prompts. $\bm{x}_{i}\in\mathbb{R}^{d}$ represents the corresponding images associated with each random string. The fine-tuning objective is to optimize the model parameters $\theta$ to minimize the loss function over the training data:
    $\theta^{*}=\arg\min_{\theta}\mathbb{E}_{(s,\bm{x})\sim\mathcal{D}_{train}}\left[\mathcal{L}(p_{\theta}(\bm{x}\mid s),\bm{x})\right]$ (19)
    where $\mathcal{L}$ denotes the loss function, such as mean squared error (MSE) or cross-entropy loss. The use of random strings as surrogate conditions aims to induce memorization within the model, causing it to overfit specific image-text associations during the fine-tuning process.
  • Attacker’s Access: After the fine-tuning phase, the attacker gains access to the fine-tuned model $p_{\theta}(\bm{x}\mid s)$. Such access enables the attacker to perform Input Manipulation. The attacker can input any text prompt $s\in\mathcal{S}_{s}$ into the model to generate corresponding images:
    $\bm{x}\sim p_{\theta}(\bm{x}\mid s)$ (20)
  • Data Extraction: By systematically querying the model with the same set of random strings $\mathcal{S}_{s}$, the attacker can generate multiple images $\{\bm{x}_{i,j}\}_{j=1}^{N_{G}}$ for each string $s_{i}$. Our goal is to identify images that closely resemble the memorized training samples:
    $\bm{x}_{i,j}\sim p_{\theta}(\bm{x}\mid s_{i}),\quad j=1,\dots,N_{G}$ (21)
  • Memorization Detection: The attacker analyzes the consistency and similarity among the generated images for each random string $s_{i}$. A low variance $\sigma_{s_{i}}^{2}$ among the samples indicates potential memorization:
    $\sigma_{s_{i}}^{2}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i,j}-\overline{\bm{x}}_{s_{i}}\|^{2}$ (22)
    where $\overline{\bm{x}}_{s_{i}}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i,j}$ is the mean image for string $s_{i}$. If $\sigma_{s_{i}}^{2}$ falls below a predefined threshold $\tau$, the attacker considers $\overline{\bm{x}}_{s_{i}}$ as a potential memorized training sample.
  • Assumptions: Knowledge of Surrogate Strings: The attacker is aware of the set of random strings $\mathcal{S}_{s}$ used during the victim’s fine-tuning process. Model Query Capability: The attacker can perform a large number of queries to the fine-tuned model with each random string $s_{i}$ to generate sufficient samples for variance analysis. No Access to Training Data: The attacker does not have direct access to the victim’s training dataset $\mathcal{D}_{train}$.
  • Attack Goals: Data Reconstruction: To reconstruct or approximate the training images $\bm{x}_{i}$ by exploiting the model’s memorization of the random string-image associations. Privacy Breach: To demonstrate the potential privacy risks associated with fine-tuned diffusion models by extracting sensitive or proprietary training data.
E.2 Extended SIDE Method
The Extended SIDE method leverages random strings as surrogate conditions to directly extract training images from a fine-tuned diffusion model. Unlike the original SIDE method, which relies on a classifier to guide the extraction process, this approach simplifies the attack by using random strings to induce and exploit memorization within the model.
Algorithm 2 Extended SIDE Method for Extracting Training Data Using Random Strings
1:Input:
  • Fine-tuned diffusion model $p_{\theta}(\bm{x}\mid s)$
  • Set of random strings $\mathcal{S}_{s}$
  • Number of samples per string $N_{G}$
  • Variance threshold $\tau$
2:Output: Extracted data $\mathcal{D}_{\text{extracted}}$
3:Initialize $\mathcal{D}_{\text{extracted}}\leftarrow\emptyset$
4:for each $s\in\mathcal{S}_{s}$ do
5:     Generate samples $\{\bm{x}_{i}^{(j)}\}_{j=1}^{N_{G}}$ where $\bm{x}_{i}^{(j)}\sim p_{\theta}(\bm{x}\mid s)$
6:     Compute mean image:
$\overline{\bm{x}}_{s}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i}^{(j)}$ (23)
7:     Compute sample variance:
$\sigma_{s}^{2}=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i}^{(j)}-\overline{\bm{x}}_{s}\|^{2}$ (24)
8:     if $\sigma_{s}^{2}<\tau$ then
9:         $\mathcal{D}_{\text{extracted}}\leftarrow\mathcal{D}_{\text{extracted}}\cup\{\overline{\bm{x}}_{s}\}$
10:     end if
11:end for
12:return $\mathcal{D}_{\text{extracted}}$
The Extended SIDE method is based on the principle that random strings, when used as surrogate conditions during fine-tuning, become strongly associated with specific training images. This follows from the results of our theoretical understanding in Theorem 1. The attacker exploits this behavior to extract training data by performing the following steps:
  1. 1.
    Selection of Random Strings: The attacker identifies or generates a set of random strings $\mathcal{S}_{s}$ that were used during the victim’s fine-tuning process. These strings act as unique keys that the model associates with specific images.
  2. 2.
    Image Generation: For each random string $s\in\mathcal{S}_{s}$, the attacker generates $N_{G}$ images by querying the fine-tuned model:
    $\bm{x}_{i}^{(j)}\sim p_{\theta}(\bm{x}\mid s),\quad j=1,\dots,N_{G}$ (25)
    The repetition of image generation for the same string allows the attacker to assess the consistency of the outputs.
  3. 3.
    Consistency Analysis: The attacker computes the mean image $\overline{\bm{x}}_{s}$ and the sample variance $\sigma_{s}^{2}$ for the generated images associated with each string $s$:
    $\overline{\bm{x}}_{s}$ $=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\bm{x}_{i}^{(j)}$ (26)
    $\sigma_{s}^{2}$ $=\frac{1}{N_{G}}\sum_{j=1}^{N_{G}}\|\bm{x}_{i}^{(j)}-\overline{\bm{x}}_{s}\|^{2}$ (27)
    A low variance $\sigma_{s}^{2}<\tau$ indicates that the generated images are highly similar, suggesting that the model has memorized the association between the string $s$ and a specific training image $\bm{x}_{s}$.
  4. 4.
    Data Extraction: If the variance $\sigma_{s}^{2}$ is below the threshold $\tau$, the attacker considers $\overline{\bm{x}}_{s}$ as an extracted training image and adds it to the dataset $\mathcal{D}_{\text{extracted}}$:
    $\mathcal{D}_{\text{extracted}}\leftarrow\mathcal{D}_{\text{extracted}}\cup\{\overline{\bm{x}}_{s}\}$ (28)
This method effectively leverages the deterministic mapping between random strings and training images induced by memorization, allowing the attacker to reconstruct the training dataset without direct access.
E.3 Algorithmic Implementation
The Extended SIDE method is formalized in Algorithm 2. The algorithm outlines the steps the attacker takes to generate images, compute the variance, and extract the training data based on the consistency of generated images. Since the attacker cannot compute similarity to the training data, they detect memorization by analyzing the diversity of generated images conditioned on the same random string $s$. The attacker identifies memorization by detecting low variance in the images generated from the same random string $s$. A low variance $\sigma_{s}^{2}$ indicates that the model consistently generates similar images for $s$, suggesting that it has memorized specific training data associated with $s$. The Extended SIDE is shown in Algorithm 2
E.4 Experiments
To verify the effectiveness of the Extended SIDE, we conduct experiments on the subset of the Laion-2B dataset, finetuning the Stable Diffusion v1.5 using the LORA[24] method. We use the following metrics to evaluate the attack results: mean SSCD (M-SSCD), AMS (mid-similarity) and LIPIPS. The results are shown in Table 5. We use the generated images of original prompts as the Baseline method. The extracted examples are shown in Figure 8.
Table 5: Performance Comparison for Extended SIDE method
Dataset Method M-SSCD AMS LIPIPS
Subset of Laion-2B Baseline 0.028 0.000 0.892
Extended SIDE(ours) 0.467 0.672 0.809
As presented in Table 5, the Extended SIDE method outperforms the baseline across all evaluation metrics:
  • M-SSCD: The SIDE method achieves an M-SSCD of 0.467 compared to the baseline’s 0.028. This significant increase indicates that SIDE extracts images with higher structural similarity to the training data, demonstrating effective memorization.
  • AMS: SIDE records an AMS of 0.672, markedly higher than the baseline’s 0.000. This suggests that SIDE successfully captures a substantial number of moderately similar images, enhancing the diversity and relevance of the extracted data.
  • LIPIPS: SIDE exhibits a lower LIPIPS of 0.809 compared to the baseline’s 0.892. The difference is less pronounced. This is because the primary purpose of LIPIPS is to evaluate perceptual differences between images, rather than structural or training-specific similarity. It measures how humans perceive image similarity based on deep features from pre-trained networks, making it effective for detecting perceptual distortions or quality changes but not for identifying whether an image was directly memorized from the training data.
Visual examples Figure 8 showcases examples of extracted images using the SIDE method alongside the corresponding training images. The visual comparison underscores the high degree of similarity achieved by SIDE, affirming its efficacy in extracting memorized training data.
Figure 8: Examples of Extracted Images using SIDE Method. Top Row: Corresponding training images from the Laion-2B subset. Bottom Row: Images generated by querying the fine-tuned Stable Diffusion model with random strings using the SIDE method. The high visual similarity demonstrates the method’s effectiveness in extracting training data.
The Extended SIDE method effectively extracts training data from fine-tuned DPMs by exploiting the memorization induced by random strings during the fine-tuning process. The quantitative results demonstrate that SIDE significantly enhances the extracted images’ structural and mid-range similarity compared to the baseline. This indicates that SIDE can reliably reproduce specific training images by leveraging the unique associations between random strings and their corresponding images.